### v1

# Generalist World Model Pre-Training for Efficient Reinforcement Learning

Abstract
Sample-efficient robot learning is a longstanding goal in robotics. Inspired by the success of scaling in vision and language, the robotics community is now investigating large-scale offline datasets for robot learning. However, existing methods often require expert and/or reward-labeled task-specific data, which can be costly and limit their application in practice. In this paper, we consider a more realistic setting where the offline data consists of reward-free and non-expert multi-embodiment offline data. We show that generalist world model pre-training (WPT), together with retrieval-based experience rehearsal and execution guidance, enables efficient reinforcement learning (RL) and fast task adaptation with such non-curated data. In experiments over 72 visuomotor tasks, spanning 6 different embodiments, covering hard exploration, complex dynamics, and various visual properties, WPT achieves 35.65% and 35% higher aggregated score compared to widely used learning-from-scratch baselines, respectively.
Reinforcement Learning, World Models, Model-based Reinforcement Learning, Continuous Control, Machine Learning
1 Introduction
Figure 1: Aggregate performance on 50 manipulation tasks from Meta-World and 22 locomotion tasks from DMControl with pixel inputs. We pre-train one single world model per benchmark on non-curated offline data, which can be fine-tuned to largely improve RL’s sample efficiency. With 150k online samples, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 using the same samples and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World).
Figure 2: Overview of generalist world model pre-training (WPT). Facing non-curated offline datasets with reward-free, non-expert, and multi-embodiment data, we train a task and embodiment-agnostic world model. The pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
With the success of scaling laws in vision and language (Brown et al., 2020; He et al., 2022; Kirillov et al., 2023; Touvron et al., 2023), the robotics community has started investigating the use of large-scale offline datasets for robot learning (Brohan et al., 2023b; Walke et al., 2023; Brohan et al., 2023a; O’Neill et al., 2023; Khazatsky et al., 2024; Team et al., 2024). Large-scale offline datasets enable the trained agents to learn a wide range of skills as well as help the agent adapt to new tasks with limited samples (Team et al., 2024).
To leverage offline datasets for decision-making, imitation learning (IL) (Brohan et al., 2023b; Team et al., 2024; Doshi et al., 2024) and offline RL (Levine et al., 2020; Kumar et al., 2020; Fujimoto & Gu, 2021; Kumar et al., 2023) are commonly used. However, these methods have limitations in terms of both data and algorithmic aspects. IL assumes the availability of expert data from human demonstration or specialist RL agents, which usually involves an expensive data collection procedure. Further to this, agents learned by IL are usually sensitive to environmental perturbations such as action delay, lighting conditions, sensor noise, etc (Chae et al., 2022; Zare et al., 2024). Offline RL aims to learn agents from offline data but suffers from training instability, especially for pixel-based observations (Kumar et al., 2020; Lu et al., 2023). Besides the training stability issue, offline RL requires task-specific datasets, labeled with rewards. This poses another challenge for many real-world applications, e.g., robotic manipulation, to utilize offline datasets for new tasks, it is necessary to retrospectively annotate the image-based offline data with rewards, which can be challenging and laborious. Furthermore, fine-tuning an offline RL agent, so-called offline-to-online RL (Nair et al., 2020; Lee et al., 2022; Zhao et al., 2022; Yu & Zhang, 2023; Nakamoto et al., 2024), requires additional tricks to stabilize training and prevent performance collapse caused by the distributional shift in the fine-tuning stage.
In contrast to IL and offline RL, which require curated datasets, many real-world scenarios have data from various sources. To this end, we focus on the more realistic setting where the offline dataset includes non-curated mixed-quality data consisting of reward-free and non-expert data, collected by multiple agents with varying embodiments. This setting has a minimal requirement for offline data, which significantly expands the pool of usable data. The primary research question we aim to answer is:
What is the best way to leverage non-curated offline data for efficient multi-task & multi-embodiment learning?
Definition 1.1 (Non-curated data).
Non-curated data encapsulates all observation-action trajectories, i.e., it can consist of reward-free and non-expert data for any embodiment.
Existing methods for leveraging offline data (e.g. IL and offline RL) typically fail given non-curated data. For example, IL would require manually filtering the dataset to select expert trajectories and offline RL would require retrospective reward labeling. With such non-curated data, a common strategy is to pre-train a visual encoder (Schwarzer et al., 2021; Nair et al., 2022; Parisi et al., 2022; Xiao et al., 2022; Yang & Nachum, 2021; Shang et al., 2024). However, visual pre-training alone fails to fully leverage the rich information that exists in the offline dataset, e.g., dynamics model, informative states, or action prior, etc.
Facing the non-curated offline dataset, training a world model seems a “natural” choice, as it holds the promise of better generalization and better sample efficiency. However, model-based approaches are not commonly used in the offline-to-online setting (Rafailov et al., 2023; Wu et al., 2024). This is partially due to the performance margin between model-based and model-free approaches not being significant (Yu et al., 2020; Kidambi et al., 2020) while model-based approaches involve higher training complexity. In addition, in the typical offline-to-online RL setting, training an offline agent with model-based approaches can be unstable, especially for pixel inputs (Lu et al., 2023), and strong regularization is required during the online training stage (Rafailov et al., 2023).
In this paper, we show that with careful design choices, world model pre-training unlocks more useable offline data and better fits the pre-train and fine-tune paradigm used for scaling. Our key insights are i) instead of pre-training a policy using model-based offline RL as done in previous work (Yu et al., 2020; Kidambi et al., 2020; Rafailov et al., 2023), learning a world model is stable and scalable, ii) unlike previous methods (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) which learns a world model per task, by simply padding actions to unify the action spaces, a single world model can be pre-trained on the non-curated offline data consisting of data sources from different tasks and different embodiments, iii) via fine-tuning, a pre-trained generalist world model can boost RL’s sample efficiency over a wide range of tasks and embodiments, iv) during fine-tuning, experience rehearsal, and execution guidance can largely improve the performance.
Table 1: Comparison with different policy learning methods that leverage offline data.
Offline RL Off2On RL RLPD MT IL MT Offline RL WPT (ours)
Reward-free offline data
Non-expert offline data
X-embodiment offline data
Continual Improvement
Training stability
We extensively evaluate WPT on 72 pixel-based continuous control tasks covering locomotion and manipulation with different action spaces, hard exploration, high dimensions, and complex dynamics. Under a limited sample budget (150k samples), WPT outperforms previous SOTA methods by a decent margin. Specifically, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 under the same sample budget and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World.). For example, WPT enables an agent learning to control an Ant robot to walk forward within 100 trails, while widely used learning-from-scratch baselines need 10-30$\times$ samples. We further demonstrate that, without any modifications, WPT also helps task adaptation, where the agent is required to continually adapt its skill to new tasks. Our results demonstrate strong performance with a large margin over model-free methods when leveraging non-curated offline data. We hope that our results motivate further investigation of model-based approaches for using offline data. In summary, our contributions include:
  • We propose a new and more realistic setting for leveraging offline data where the offline data consists only of reward-free and non-expert multi-embodiment data.
  • We show that generalist world model pre-training along with a series of careful design choices achieves strong performance in 72 visuomotor tasks spanning 6 embodiments.
  • To facilitate further research with non-curated offline data, we open-source our datasets, including 60K trajectories and 10M transitions from the DMControl and Meta-World benchmarks.
2 Related Work
In this section, we review methods which leverage offline data, including pre-training in the context of RL and the so-called generalist agents. See Table 1 for a comparison of what types of data the methods can use.
RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve the sample efficiency in RL. One representative method is Offline RL. Offline RL trains agents from offline data without interacting with the environment. It usually constrains the distance between the learned policy and behavior policies in different ways (Kumar et al., 2019, 2020; Wu et al., 2019; Kostrikov et al., 2021, 2022; Fujimoto & Gu, 2021; Uchendu et al., 2023). However, the performance of the learned policy is highly affected by the quality of offline datasets (Yarats et al., 2022). To continue improving the policy, offline-to-online RL (Nair et al., 2020; Zhao et al., 2022; Lee et al., 2022; Yu & Zhang, 2023; Rafailov et al., 2023) was proposed.
Offline RL and offline-to-online RL commonly suffer from training instability and require additional treatments (Lee et al., 2022; Lu et al., 2023). RLPD (Ball et al., 2023) shows that off-policy RL delivers strong performance by directly concatenating the offline data with online data. However, RLPD still requires reward-labeled task-specific offline data and hasn’t discussed the multi-embodiment cases. Recently, ExPLORe (Li et al., 2023) labels reward-free offline data with approximated upper confidence bound (UCB) of rewards to solve hard exploration tasks. However, ExPLORe obtains good performance when the offline data includes near-expert data of the target tasks, while we consider a more general setting with non-curated offline data.
RL with multi-task offline datasets   Recently, several works have applied offline RL in a multi-task setting (Kalashnikov et al., 2021; Yu et al., 2021; Kumar et al., 2023; Julian et al., 2020; Hansen et al., 2024), but require known rewards.  Georgiev et al. (2024) pre-trains a world model for multi-task RL but is designed for state-based inputs and requires reward-labeled offline data for world model pre-training.
To lift the assumption of known rewards, human labeling (Cabi et al., 2020; Singh et al., 2019), inverse RL (Ng et al., 2000; Abbeel & Ng, 2004) or generative adversarial imitation learning (Ho & Ermon, 2016) can be applied. However, this requires human labor or expert demonstrations. Yu et al. (2022) directly sets zero rewards for unlabelled data, which introduces additional bias.
Another stream of work leverages in-the-wild data for RL training. Most methods focus on representation learning. Stooke et al. (2021); Yang & Nachum (2021); Schwarzer et al. (2021); Shah & Kumar (2021); Yuan et al. (2022); Wang et al. (2022); Parisi et al. (2022); Sun et al. (2023); Ze et al. (2023) investigate different representation learning methods and show promising performance boosts during fine-tuning.
Generalist Agents   RL methods usually perform well on a single task (Vinyals et al., 2019; Andrychowicz et al., 2020), however, this contrasts with humans that can perform multiple tasks well. Recent works have proposed generalist agents that master a diverse set of tasks with a single agent (Reed et al., 2022; Brohan et al., 2023b; Team et al., 2024). These methods typically resort to scalable models and large datasets and are trained via imitation learning (Brohan et al., 2023b, a; O’Neill et al., 2023; Khazatsky et al., 2024). In contrast, we seek to train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
World models   World models learn to predict future observations or states based on historical information. World models have been widely investigated in online model-based RL (Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024). Recently, the community has started investigating scaling world models (Ha & Schmidhuber, 2018), for example, Hu et al. (2023a); Pearce et al. (2024); Wu et al. (2025); Agarwal et al. (2025) train world model with Diffusion Models or Transformers. However, these models are usually trained on demonstration data. In contrast, we explore the offline-to-online RL setting – closely fitting the successful pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and non-expert data to increase the amount of available data for pre-training.
3 Methods
In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a generalist (i.e. multi-task & embodiment) world model, given offline data, which rather importantly, includes reward-free and non-expert data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion. See Fig. 2 and Alg. 1.
3.1 Problem Setup
In this paper, we assume the agent can access a static offline dataset $\mathcal{D}_{\text{off}}$, which consists of trajectories $\{\tau_{\text{off}}^{i}\}_{i=1}^{N_{\text{off}}}$, where $\tau_{\text{off}}^{i}$ includes observations and actions $\{o^{i}_{t},a^{i}_{t}\}_{t=0}^{T}$ collected by unknown behavior policies. This means that (i) rewards $r^{i}_{t}$ are unknown, (ii) $\mathcal{D}_{\text{off}}$ does not necessarily include expert trajectories, and (iii) datasets consist of multi-embodiment data. The agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$. The goal of this paper is to learn a high-performance policy from both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ whilst minimizing the amount of online interaction ($N_{on}$) by leveraging the unlabeled offline data $\mathcal{D}_{\text{off}}$. Removing the requirement of rewards and expert trajectories enables the agent to leverage a large set of diverse datasets.
3.2 Multi-Embodiment World Model Pre-training
During pre-training, instead of training one model per task – as done in previous work (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) – we show that a single multi-task & embodiment world model more effectively leverages the offline data. Compared to methods (Yang & Nachum, 2021; Schwarzer et al., 2021; Yuan et al., 2022; Parisi et al., 2022; Ze et al., 2023) that only pre-train representations, pre-training a world model creates a general understanding of the environment, which we show can further boost agents’ performance during fine-tuning.
We adopt a widely used world model in the literature, the recurrent state space model (RSSM) (Hafner et al., 2019) by making several modifications: (i) we remove the task-related losses, (ii) we pad the action with zeros to unify the action dimension of different embodiments, and (iii) we scale the model size to 280M. Given these modifications, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for mastering different tasks. Our first stage consists of pre-training the following components:
Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})$
Encoder $:z_{t}\sim q_{\theta}(z_{t}|h_{t},o_{t})$
Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}|h_{t})$
Decoder $:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}|h_{t},z_{t}).$
The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are optimized jointly by minimizing:
$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}|z_{t},h_{t})}_{\text{pixel reconstruction loss}}\\ +\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}|h_{t},o_{t})\ ||\ p_{\theta}(z_{t}|h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
The first term minimizes the reconstruction error while the second term enables the latent dynamics learning. Note that we removed the task-related objectives, e.g., reward prediction and continue prediction. We note that there is plenty of room to improve the world model’s pre-training, e.g., by leveraging recently developed self-supervised training methods (Eysenbach et al., 2023), or advanced architectures (Vaswani et al., 2017; Gu et al., 2022). These improvements are orthogonal to our method and we leave it as future work.
3.3 RL-based Fine-tuning
In our fine-tuning stage, we allow the agent to interact with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$. The newly collected data is used to learn a reward function via supervised learning whilst also fine-tuning the world model with Eq. 1. For simplicity, we represent the concatenate of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$. The actor and critic are trained with imagined trajectories $\tilde{\tau}(s,a)$ generated by rolling out the policy $\pi_{\phi}(a|s)$ in the world model $p_{\theta}$ starting from the initial state distribution $p_{0}(s)$, which are samples from replay buffer. Following DreamerV3, the critic $v_{\phi}(V_{t}^{\lambda}\mid s_{t})$ learns to approximate the distribution
over the $\lambda$-return $V_{t}^{\lambda}$ calculated as,
$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases},$ (2)
where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\mid s_{t})]$ denotes the expected value of the distribution predicted by the critic. It is trained using maximum likelihood by minimizing
$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\mid s_{t})\right].$ (3)
The actor $\pi_{\phi}(a_{t}|s_{t})$ is then updated by maximizing the $\lambda$-return regularized by actor entropy $\mathbf{H}[a_{t}|s_{t}]$:
$\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}|s_{t}]\right)\right].$ (4)
In practice, we found that using the pre-trained world model alone fails to work well in many cases, especially for hard-exploration tasks. To identify the reasons, we need to rethink the role of each component in policy updates. Given the policy update rule in Eq. 4, the policy update is infected by the distribution of imagined trajectories $\tilde{\tau}(s,a)=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}|s_{t})p_{\theta}(s_{t+1}|s_{t},a_{t})$ and the reward model $r_{\theta}(s_{t},a_{t})$, i.e., the initial state distribution $p_{0}(s)$, the policy $\pi_{\phi}(a|s)$, the dynamics model $p_{\theta}(s_{t+1}|s_{t},a_{t})$, and the reward model $r_{\theta}(s_{t},a_{t})$. Intuitively, as shown in Fig. 3, the policy will learn a good policy if the imagined trajectories $\tilde{\tau}$ contain “promising” data for policy learning and the reward function $r_{\theta}(s_{t},a_{t})$ can correctly label those trajectories. The world model pre-training gives a better initialization of $p_{\theta}(s_{t+1}|s_{t},a_{t})$. However, when the online data distribution is narrow – as in hard exploration tasks – the initial state distribution $p_{0}(s)$ can also be narrow, which prevents the actor from reaching high-reward states during imagination. Also, the reward function $r_{\theta}(s,a)$ trained on the narrow online data distribution will incorrectly assign rewards for transitions generated by the world model pre-trained on a broader distribution. That is, the agent fails to improve the policy even if “promising” trajectories exist in $\tilde{\tau}$, since incorrect rewards are assigned. Furthermore, fine-tuning the world model $p_{\theta}(s_{t+1}|s_{t},a_{t})$ on the narrow distribution causes catastrophic forgetting. In the following, we improve the agents’ performance by reusing non-curated offline data for i) augmenting the initial state distribution $p_{0}(s)$, ii) collecting diverse online trajectories, and iii) improving the reward predictions. See Fig. 3 for an illustrative motivation.
Retrieval-based Experience Rehearsal
Figure 3: Motivation of using experience rehearsal and execution guidance. During fine-tuning, experience rehearsal augments the initial state $p_{0}(s)$ used for model rollout, which enables the model rollout to start from informative states. Execution guidance helps the agent to collect data that is close to the offline data distribution, which enriches the diversity of the collected data and improves the reward model’s accuracy on a broader state distribution.
Experience replay is a simple yet effective approach to prevent catastrophic forgetting (Khetarpal et al., 2022) as well as improving policy learning (Ball et al., 2023; Li et al., 2023). In the RL setting, Ball et al. (2023) demonstrates promising results by replaying offline data to improve sample efficiency but requires the offline data to be reward-labeled. ExPLORe (Li et al., 2023) reuses single-embodiment but reward-free offline data to encourage exploration by labeling offline data with UCB rewards. Unlike the replay scheme used in ExPLORe, where the offline datasets are relatively small and well-structured, directly replaying the non-curated offline data can be challenging. Our dataset can be 100$\sim$ 100$\times$ 100 × larger and consist of different tasks and embodiments. As such, directly labeling the offline dataset with UCB reward or even replaying the dataset can be infeasible.
To enable the usage of the non-curated offline data, we propose a simple retrieval-based approach to filter the trajectories that are close to the downstream tasks. Specifically, we retrieve a subset of trajectories from the non-curated offline data based on the neural feature distance between the online samples and the trajectories in the offline dataset. We calculate the distance as :
$\mathbf{D}=||\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})||_{2},$ (5)
where $\text{e}_{\theta}$ is a feature extractor and $o_{\text{on}}$ and $o_{\text{off}}$ are images from the initial observation of trajectories from the online buffer and the offline dataset, respectively. In practice, we found that neural representation matters for retrieval and the encoder learned via world model pre-training works better than the general visual model R3M (Nair et al., 2022).
To efficiently search over the whole dataset, we pre-build the key-value pairs between the trajectories ID and its neural features and use Faiss (Douze et al., 2024) for similarity search. Once the key-value database is built, the buffer retrieval procedure only takes seconds. The retrieved offline buffer, although reward-free, contains embodiment-specific data and is feasible to use. The retrieved offline data can be used to prevent catastrophic forgetting in the world model and to augment the initial state distribution $p_{0}(s)$ used for trajectory imagination.
Figure 4: WPT enables sample-efficient RL training. We include 6 tasks from DMControl and Meta-World with pixel inputs. The mean and the corresponding 95% confidence interval across 3 seeds are plotted. Full results on 72 tasks are included in App. D.
Execution Guidance via Prior Actors
The common practice in RL training is to initialize the replay buffer with a random policy and gradually collect new data by interacting with the environment using the training policy. When using a pre-trained world model, it is preferred to start the training from a distribution that is close to the offline data distribution. This has three benefits: First, offline data usually contains useful information for policy training, such as near-expert trajectories or a broader state-action distribution. Second, the distribution shift between the offline and online data may ruin the pre-trained weights, so bringing the initial distribution close to the offline data is a good strategy to prevent this. Thirdly, the reward function can get better predictions for states that are close to the offline data distribution. This improves policy learning as the initial states used to generate the imagined trajectories $\tilde{\tau}$ are augmented with offline samples.
To guide the online data collection such that it remains close to the offline data distribution, we simply train a prior actor $\pi_{\text{bc}}$ via behavior cloning on the retrieved buffer. During online data collection, we switch between the policy prior $\pi_{\text{bc}}$ and the RL agent $\pi_{\phi}$ according to a pre-defined schedule. Specifically, at the beginning of an episode, we decide whether to use $\pi_{\text{bc}}$ according to the scheduled probability. If the $\pi_{\text{bc}}$ is selected to use, we then randomly select the starting time step $t_{\text{bc}}$ and duration $H$ of using $\pi_{\text{bc}}$. $\pi_{\phi}$ is then used for the rest of the time steps. This procedure is similar to JSRL (Uchendu et al., 2023) but with a couple of differences: i) WPT leverages non-curated offline data while JSRL uses task-specific offline data and ii) WPT shows the superiority of a model-based approach for leveraging offline data while JSRL is a model-free approach. As a result, we show that WPT obtains better performance than JSRL on tested tasks. The full algorithm is presented in App. C.
4 Experiments
In this section, we investigate our WPT in two different RL settings. In the first setting, we focus on how the proposed WPT helps to improve the sample efficiency of RL training. In the second setting, we investigate whether WPT enables fast task adaptation, which is particularly important to empower the agent’s life-long learning ability. Lastly, we conduct an ablation study to investigate the role of proposed design choices.
4.1 Dataset
Our dataset consists of data from two different benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig. 2 and Fig. 9 for the full list. For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents (Pathak et al., 2017; Rajeswar et al., 2023), which are trained via curiosity without any task-related information. For Meta-World, we collect 50k trajectories spinning 50 tasks by executing the pre-trained RL agents from TDMPCv2 (Hansen et al., 2024) with injected Gaussian noise with $\sigma=[0.1,0.3,0.5,1.0,2.0]$ to mimic imperfect demonstrations. In total, the offline dataset consists of 60k trajectories with 10M state-action pairs from 6 different embodiments.
4.2 WPT Enables Sample-efficient Fine-tuning
Tasks   We evaluate our method on six selected pixel-based continuous control tasks from DMControl and Meta-World. The chosen tasks cover different challenges in RL. Except for learning from high-dimensional observations, we include a hard exploration version of Cheetah Run by setting zero rewards below a reward threshold and applying an action penalty. We include Walker Run and Quadruped Walk for complex dynamics. We further include challenging manipulation tasks from MetaWorld.
Baselines   In our experiments, we consider four different baselines for leveraging reward-free offline data. Note that as these baselines are not designed to handle multi-embodiment offline data, we preprocess the offline data to only contain task-related data for the baselines. Although this is an unfair comparison for WPT as baselines eliminate the difficulty raised with the large non-curated dataset, we still outperform these baselines by a large margin on tested tasks. The compared baselines are:
  • R3M pre-trains visual representation aiming to boost policy learning. We use a ResNet-18 backbone and train additional policy and value heads via RL. We include this baseline to compare with a representative approach for visual pre-training.
  • UDS suggests to label offline data with zero. UDS is an offline RL method, which cannot be directly used for comparison. We instead set zero reward for offline data and use RLPD (Ball et al., 2023) for policy training.
  • ExPLORe labels offline data with UCB reward. We follow the implementation from Li et al. (2023) but use ensembles for reward estimation, which has better performance.
  • JSRL-BC (Uchendu et al., 2023) collects online data with both the training policy and a prior policy trained on offline data via IL. We include JSRL-BC to compare with different ways of using offline data.
Figure 5: WPT enables fast task adaptation. We train an RL agent to control an Ant robot from DMControl to finish a series of tasks incrementally. WPT outperforms the widely used baseline by a decent margin by properly leveraging non-curated offline data.
Figure 6: Ablation study on the role of each component. “P” represents world model pretraining, “ER” means experience rehearsal, and “G” represents execution guidance. Together with the proposed retrieval-based experience rehearsal and execution guidance, world model pre-training boosts RL performance on a wide range of tasks.
Results   Figure 4 shows comparison results with baselines. Our method outperforms or matches all compared baselines by a large margin. Compared to R3M, WPT stresses the importance of world model pre-training as well as reusing offline data during fine-tuning compared to representation learning only. R3M fails to improve sample efficiency on most of the tested tasks. The same conclusion is obtained in Hansen et al. (2023). UDS and ExPLORe reuse offline data by labeling offline data with zero rewards and UCB rewards respectively. The labeled offline data is concatenated with online data for off-policy updates. We find UDS only shows slightly better performance on the Walker Run task compared to R3M and JSRL-BC, showing the ineffectiveness of directly labeling offline data with zero rewards. ExPLORe shows better performance on 2/3 locomotion tasks compared to other baselines and shows signs of meaningful progress on two challenging manipulation tasks. However, WPT still outperforms ExPLORe by a large margin. Compared to ExPLORe, WPT demonstrates the superiority of world model pre-training as well as execution guidance.
Compared to JSRL-BC, our method also demonstrates a clear performance boost. JSRL-BC’s performance heavily depends on the distribution of offline datasets. It can achieve good performance if a good prior actor can be extracted from the offline data. However, in our case, since we aim to lift the assumption of expert trajectories, the pre-trained BC agent usually fails to work well in the target task. As a result, JSRL-BC is only slightly better than the rest baselines on the Cheetah Run Hard, Assembly, and Shelf Place tasks while achieving similar results on the rest. In contrast, WPT works well on leveraging non-expert offline data. For example, the Quadruped Walk shows that WPT can nicely benefit from exploratory offline data collected by unsupervised RL, which enables it to control the Ant robot to walk forward from pixel inputs with only 100 trials.
In Fig. 1 and App. D, we further compare our method with two widely used learning-from-scratch baselines, Dreamerv3 and DrQv2, on 72 tasks. We show that by leveraging the non-curated offline data, WPT can clearly improve the sample efficiency on a wide range of tasks. Furthermore, our method achieves promising performance on several hard exploration tasks, while learning-from-scratch baselines fail.
4.3 WPT Enables Fast Task Adaptation
In many real-world applications, we want the robot to continuously adapt to new tasks. In this section, we investigate whether WPT benefits task adaptation. In this setting, the agent is required to incrementally solve a sequence of tasks. The setting is very similar to continual reinforcement learning (CRL) or life-long RL (Parisi et al., 2019; Khetarpal et al., 2022) but with a limited set of tasks for simplicity. Note that CRL has a broad scope; assumptions and experiment setups vary among methods, which makes it difficult to set up a fair comparison with other methods. Instead of proposing a SOTA CRL method, we aim to demonstrate that our method offers a simple yet general recipe to leverage previous data that also fits the CRL setting.
Setup & baselines   We set our continual multi-task adaption experiment based on the Ant robot from the DeepMind Control Suite. Specifically, the agent needs to sequentially solve stand, walk, run, jump, roll, and roll fast tasks. We train 300K environment steps for each task. To solve one task, the agent accesses all previous experiences as well as the model’s weights trained on previous tasks. To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain. We compare our method with PackNet (Mallya & Lazebnik, 2018), which is a common baseline for continual learning. PackNet iteratively prunes the actor’s parameters by keeping the weights with a larger magnitude while re-initializing the rest. By doing so, the actor maintains skills learned previously, which may help the learning of new tasks. For each new task, we fine-tune the actor model by iterative pruning while randomly initializing the critic model as rewards are not shared among tasks.
Results   Figure 5 shows policy performance on incremental task adaption. As we can see WPT outperforms baselines by a large margin. Compared to PackNet, WPT shows great potential for pre-training of world models with non-curated offline data together with experience rehearsal and actor prior. For the tested Ant domain, WPT enables the agent to adapt to new tasks within only 100 trials on all six tasks. With the limited sample budget, PackNet only achieves $\sim 20-60\%$ episodic returns of WPT. We would argue the diversity of the non-curated offline data contributes to WPT’s supreme performance on task adaptation.
4.4 Ablations
In this section, we investigate the role of each component of our method. We use the same set of tasks used in Fig. 4. As shown in Fig. 6, world model pre-training shows promising results when the offline dataset consists of diverse trajectories, such as data collected by exploratory agents, while fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks. This is due to i) world model pre-training alone failing to fully leverage the rich action prior in the offline data and ii) the distributional shift between offline and online data hurts the world model fine-tuning. The proposed retrieval-based experience rehearsal and execution guidance helps the agent benefit from action prior in the offline data and enriches the online sample diversity, which together enables WPT to achieve strong performance on a wide range of tasks.
5 Conclusion
We propose WPT, a simple yet efficient approach to leverage non-curated offline data. We show that the generalist world model pre-trained on non-curated data can boost RL training spanning multiple tasks and multiple embodiments. Together with retrieval-based experience rehearsal and execution guidance, WPT outperforms baselines on a wide range of tasks, including 22 locomotion tasks and 50 manipulation tasks. WPT unlocks ample sources of offline data and our results show that leveraging this non-curated data leads to strong performance. Nevertheless, our method can be improved in multiple ways, for example, by extending our methods to real-world applications as well as proposing novel world model architectures.
Impact Statement
This paper contributes to the field of reinforcement learning (RL), with potential applications including robotics and autonomous machines. While our methods hold promise for advancing technology, they could also be applied in ways that raise ethical concerns, such as in autonomous machines exploring the world and making decisions on their own. However, the specific societal impacts of our work are broad and varied, and we believe a detailed discussion of potential negative uses is beyond the scope of this paper. We encourage a broader dialogue on the ethical use of RL technology and its regulation to prevent misuse.
Appendices
Appendix A Implementation Details
A.1 Behavior Cloning
The Behavior Cloning methods used in both the execution guidance of WPT and JSRL-BC are the same. We use a four-layer convolutional neural network (LeCun et al., 1995) with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm (Ba et al., 2016) after all linear layers.
We list the used encoder and actor architectures for reference.
1class Encoder(nn.Module):
2 def __init__(self, obs_shape):
3 super().__init__()
4 assert obs_shape == (9, 64, 64), f’obs_shape is {(obs_shape)}, but expect (9, 64, 64)’ # inputs shape
5
6 self.repr_dim = (32 * 8) * 2 * 2
7 _input_channel = 9
8
9 self.convnet = nn.Sequential(
10 nn.Conv2d(_input_channel, 32, 4, stride=2), # [B, 32, 31, 31]
11 nn.ELU(),
12 nn.Conv2d(32, 32*2, 4, stride=2), #[B, 64, 14, 14]
13 nn.ELU(),
14 nn.Conv2d(32*2, 32*4, 4, stride=2), #[B, 128, 6, 6]
15 nn.ELU(),
16 nn.Conv2d(32*4, 32*8, 4, stride=2), #[B, 256, 2, 2]
17 nn.ELU())
18 self.apply(utils.weight_init)
19
20 def forward(self, obs):
21 B, C, H, W = obs.shape
22
23 obs = obs / 255.0 - 0.5
24 h = self.convnet(obs)
25 # reshape to [B, -1]
26 h = h.view(B, -1)
27 return h
1class Actor(nn.Module):
2 def __init__(self, repr_dim, action_shape, feature_dim=50, hidden_dim=1024):
3 super().__init__()
4 self.trunk = nn.Sequential(nn.Linear(repr_dim, feature_dim),
5 nn.LayerNorm(feature_dim), nn.Tanh())
6
7 self.policy = nn.Sequential(nn.Linear(feature_dim, hidden_dim),
8 nn.LayerNorm(hidden_dim), nn.ELU(),
9 nn.Linear(hidden_dim, hidden_dim),
10 nn.LayerNorm(hidden_dim), nn.ELU(),
11 nn.Linear(hidden_dim, action_shape[0]))
12
13 self.apply(utils.weight_init)
14
15 def forward(self, obs, std):
16 h = self.trunk(obs)
17 reutrn self.policy(h)
A.2 JSRL+BC
Jump-start RL (Uchendu et al., 2023) is proposed as an offline-to-online RL method. It includes two policies, a prior policy $\pi_{\theta_{1}}(a|s)$ and a behavior policy $\pi_{\theta_{2}}(a|s)$, where the prior policy is trained via offline RL methods and the behavior policy is updated during the online learning stage. However, offline RL requires the offline dataset to include rewards for the target task. To extract behavior policy from the offline dataset, we use the BC agent described above as the prior policy. During online training, in each episode, we randomly sample the rollout horizon $h$ of the prior policy from a pre-defined array np.arange(0, 101, 10). We then execute the prior policy for $h$ steps and switch to the behavior policy until the end of an episode.
A.3 ExPLORe
For the ExPLORe baseline, we follow the original training code   [^1] Source code of ExPLORe https://github.com/facebookresearch/ExPLORe. We sweep over several design choices: i) kernel size of the linear layer used in the RND and reward models: [256 (default), 512]; ii) initial temperature value: [0.1 (default), 1.0]; iii) whether to use LayerNorm Layer (no by default); iv) learning rate: [1e-4, 3e-4 (default)]. However, we fail to obtain satisfactory performance. There are several potential reasons: i) the parameters used in the ExPLORe paper are tuned specifically to their setting, where manipulation tasks and near-expert trajectories are used; ii) the coefficient term of the RND value needs to be tuned carefully for different tasks and the reward should also be properly normalized.
To achieve reasonable performance and eliminate the performance gap caused by implementation-level details, we make the following modifications: i) we replace the RND module with ensembles to calculate uncertainty; ii) the reward function share the latent space with actor and critic.
Appendix B More Ablation Study
We include per-task results used in Sec. 4.4 in Fig. 7.
Figure 7: Ablation study of the role of each component.
We further compare results using uncertainty-based rewards as suggested in (Li et al., 2023) compared with our expectation guidance. Instead of using UCB for reward labels, we use Optimistic Thompson Sampling (OTS) (Hu et al., 2023b), which is less sensitive w.r.t the selection of hyperparameters. As shown in Fig. 8, our method outperforms the variant using OTS on Assembly and Stick Pull by a large margin, showing the effectiveness of using execution guidance.
Figure 8: Ablation study of using uncertainty-based reward labeling.
Appendix C Algorithm
The full algorithm is described in Alg. 1.
Algorithm 1 Generalist World Model Pre-Training for Efficient RL
0:  Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$              World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$              Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
   // Task-Agnostic World Model Pre-Training
  for num. pre-train steps do
     Randomly sample mini-batch $\mathcal{B_{\text{off}}}:\{o_{t},a_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{\text{off}}$.
     Update world model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$ by minimizing Eq. 1 on sampled batch $\mathcal{B}$.
  end for
   // Task-Specific Training
   // Experience Retrieval
  Collect one initial observation $o^{0}_{\text{on}}$ from the environment.
  Compute the visual similarity between $o_{\text{on}}$ and initial observations of trajectories $o_{\text{off}}$ in $\mathcal{D}_{\text{off}}$ using Eq. 5.
  Select R trajectories according to Eq. 5 and fill $\mathcal{D}_{\text{retrieval}}$.
   // Behavior Cloning Policy Training
  for num. bc updates do
     Randomly sample mini-batch $\mathcal{B_{\text{retrieval}}}:\{o_{i},a_{i}\}_{i=0}^{N}$ from $\mathcal{D}_{\text{retrieval}}$.
     Update $\pi_{\phi_{\text{BC}}}$ by minimizing $-\frac{1}{N}\sum_{i=0}^{N}\log\pi_{\phi_{\text{BC}}}(a_{t}|o_{t})$.
  end for
   // Task-Specific RL Fine-Tuning
  for num. episodes do
      // Collect Data
     Decide whether to use $\pi_{\phi_{\text{BC}}}$ according to the predefined schedule.
     if Select $\pi_{\phi_{\text{BC}}}$ then
        Randomly select the starting time step $k$ and the rollout horizon $H$.
     end if
     $t\leftarrow 0$
     while  $t\leq$ episode length do
        $a_{t}=\pi_{\phi_{BC}}(a_{t}|o_{t})$ if Use $\pi_{\phi_{\text{BC}}}$ and $k\leq t\leq H$ else $a_{t}=\pi_{\phi_{RL}}(a_{t}|o_{t})$.
        Interact the environment with $a_{t}$. Store $\{o_{t},a_{t},r_{t},o_{t+1}\}$ to $\mathcal{D}_{\text{on}}$.
        $t\leftarrow t+1$
     end while
      // Update Models
     for num. grad steps do
        Randomly sample mini-batch $\mathcal{B}_{\text{on}}:\{o_{t},a_{t},r_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{on}$ and $\mathcal{B}_{\text{retrieval}}:\{o_{t},a_{t},r_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{\text{retrieval}}$.
        Update world model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$ by minimizing Eq. 1 on sampled batch $\{\mathcal{B_{\text{on}},\mathcal{B}_{\text{retrieval}}}\}$.
[…]
         // Update policy and value function
        Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with policy $\pi_{\phi_{\text{RL}}}$.
        Update policy $\pi_{\phi_{\text{RL}}}$ and value function $v_{\phi}$ with Eq. 4.
     end for
  end for
Appendix D Full Results
In Table 2 and Table 3, we list the success rate of 50 Meta-World benchmark tasks with pixel inputs. In Table 4, we list the episodic return of DMControl of 22 tasks. We compare WPT at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples. We mark the best result with a bold front at 150k samples and use underlining to mark the highest score overall.
D.1 Meta-World Benchmark
Table 2: Success rate of Meta-World benchmark with pixel inputs.
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
Assembly 0.0 0.0 0.0 0.0 0.2
Basketball 0.0 0.97 0.0 0.0 0.4
Bin Picking 0.0 0.93 0.0 0.33 0.8
Box Close 0.13 0.9 0.0 0.0 0.9
Button Press 1.0 0.7 0.47 0.13 0.9
Button Press Topdown 1.0 1.0 0.33 0.17 1.0
Button Press Topdown Wall 1.0 1.0 0.73 0.63 1.0
Button Press Wall 1.0 1.0 0.93 0.77 1.0
Coffee Button 1.0 1.0 1.0 1.0 1.0
Coffee Pull 0.6 0.8 0.0 0.6 0.6
Coffee Push 0.67 0.77 0.13 0.2 0.7
Dial Turn 0.67 0.43 0.13 0.17 0.67
Disassemble 0.0 0.0 0.0 0.0 0.0
Door Close - - - - 1.0
Door Lock 1.0 0.93 0.6 0.97 0.9
Door Open 1.0 0.97 0.0 0.0 0.8
Door Unlock 1.0 1.0 1.0 0.63 0.8
Drawer Close 0.93 1.0 0.93 1.0 0.9
Drawer Open 0.67 0.33 0.13 0.33 1.0
Faucet Open 1.0 1.0 0.47 0.33 1.0
Faucet Close 0.87 1.0 1.0 1.0 0.8
Hammer 1.0 1.0 0.07 0.4 1.0
Hand Insert 0.07 0.57 0.0 0.1 0.4
Handle Press Side 1.0 1.0 1.0 1.0 1.0
Handle Press 1.0 1.0 0.93 0.97 1.0
Handle Pull Side 0.67 1.0 0.67 0.6 1.0
Handle Pull 0.67 0.6 0.33 0.6 1.0
Lever Pull 0.73 0.83 0.0 0.33 0.8
More results in the next page, see Table 3
Table 3: Success rate of Meta-World benchmark with pixel inputs (Cont.).
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
Peg Insert Side 1.0 1.0 0.0 0.27 1.0
Peg Unplug Side 0.93 0.9 0.53 0.5 0.8
Pick Out of Hole 0.0 0.27 0.0 0.0 0.3
Pick Place Wall 0.2 0.17 0.0 0.0 0.5
Pick Place 0.67 0.67 0.0 0.0 0.2
Plate Slide Back Side 1.0 1.0 0.93 1.0 1.0
Plate Slide Back 1.0 1.0 0.8 0.97 1.0
Plate Slide Side 1.0 0.9 0.73 0.5 0.5
Plate Slide 1.0 1.0 0.93 1.0 1.0
Push Back 0.33 0.33 0.0 0.0 0.2
Push Wall 0.33 0.57 0.0 0.0 0.9
Push 0.26 0.93 0.0 0.13 0.7
Reach 0.87 0.73 0.67 0.43 0.3
Reach Wall 1.0 0.87 0.53 0.7 0.9
Shelf Place 0.4 0.43 0.0 0.0 0.87
Soccer 0.6 0.3 0.13 0.13 0.67
Stick Push 0.0 0.07 0.0 0.0 0.4
Stick Pull 0.0 0.33 0.0 0.0 0.67
Sweep Into 0.87 1.0 0.0 0.87 0.9
Sweep 0.0 0.73 0.0 0.3 0.6
Window Close 1.0 1.0 0.93 1.0 0.8
Window Open 1.0 0.97 0.6 1.0 0.9
Mean 0.900 0.753 0.360 0.442 0.750
Medium 0.870 0.900 0.130 0.330 0.800
D.2 DMControl Benchmark
Table 4: Episodic return of DMControl benchmark with pixel inputs.
Tasks Dreamer v3 @ 500k DrQ v2 @ 500k Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
CartPole Balance 994.3 992.3 955.8 983.3 995.0
Acrobot Swingup 222.1 30.3 85.2 20.8 50.3
Acrobot Swingup Sparse 2.5 1.17 1.7 1.5 26.9
Acrobot Swingup Hard -0.2 0.3 2.0 0.4 10.7
Walker Stand 965.7 947.6 946.2 742.9 969.4
Walker Walk 949.2 797.8 808.9 280.1 959.1
Walker Run 616.6 299.3 224.4 143.0 728.0
Walker Backflip 293.6 96.7 128.2 91.7 306.0
Walker Walk Backward 942.9 744.3 625.9 470.9 863.6
Walker Walk Hard -2.1 -9.5 -4.7 -17.1 878.3
Walker Run Backward 363.8 246.0 229.4 167.4 349.3
Cheetah Run 843.7 338.1 621.4 251.2 526.1
Cheetah Run Front 473.8 202.4 143.1 108.4 360.5
Cheetah Run Back 657.4 294.4 407.6 171.2 446.0
Cheetah Run Backwards 693.8 384.3 626.6 335.6 542.2
Cheetah Jump 597.0 535.6 200.8 251.8 634.1
Quadruped Walk 369.3 258.1 145.2 76.5 933.6
Quadruped Stand 746.0 442.2 227.2 318.9 936.4
Quadruped Run 328.1 296.5 183.0 102.8 802.5
Quadruped Jump 689.6 478.3 168.3 190.5 813.5
Quadruped Roll 663.9 446.0 207.9 126.2 970.8
Quadruped Roll Fast 508.8 366.9 124.8 164.7 782.0
Mean 541.81 372.23 320.86 226.49 631.10
Medium 606.8 318.70 204.35 166.05 755.0
Appendix E Hyperparameters
In this section, we list important hyperparameters used in WPT.
Table 5: Hyperparameters used in WPT.
Hyperparameter Value
Pre-training
Stacked images 1
Pretrain steps 200,000
Batch size 16
Sequence length 64
Replay buffer capacity Unlimited
Replay sampling strategy Uniform
RSSM
     Hidden dimension 12288
     Deterministic dimension 1536
     Stochastic dimension 32 * 96
     Block number 8
     Layer Norm True
CNN channels [96, 192, 384, 768]
Activation function SiLU
Optimizer
     Optimizer Adam
     Learning rate 1e-4
     Weight decay 1e-6
     Eps 1e-5
     Gradient clip 100
Fine-tuning
Warm-up frames 15000
Execution Guidance Schedule linear(1,0,50000) for DMControl
linear(1,0,1,150000) for Meta-Wolrd
Action repeat 2
Offline data mix ratio 0.25
Discount 0.99
Discount lambda 0.95
MLPs [512, 512, 512]
MLPs activation SiLU
Actor critic learning rate 8e-5
Actor entropy coef 1e-4
Target critic update fraction 0.02
Imagine horizon 16
Appendix F Task Visualization
Figure 9: Visualization of tasks from DMControl and Meta-World used in our paper.

### v2

@@
-Generalist World Model Pre-Training for Efficient Reinforcement Learning
+Efficient Reinforcement Learning by Guiding Generalist World Models with Non-Curated Data
 Abstract
-Sample-efficient robot learning is a longstanding goal in robotics.
-Inspired by the success of scaling in vision and language, the robotics community is now investigating large-scale offline datasets for robot learning.
-However, existing methods often require expert and/or reward-labeled task-specific data, which can be costly and limit their application in practice.
-In this paper, we consider a more realistic setting where the offline data consists of reward-free and non-expert multi-embodiment offline data.
-We show that generalist world model pre-training (WPT), together with retrieval-based experience rehearsal and execution guidance, enables efficient reinforcement learning (RL) and fast task adaptation with such non-curated data.
-In experiments over 72 visuomotor tasks, spanning 6 different embodiments, covering hard exploration, complex dynamics, and various visual properties, WPT achieves 35.65% and 35% higher aggregated score compared to widely used learning-from-scratch baselines, respectively.
-Reinforcement Learning, World Models, Model-based Reinforcement Learning, Continuous Control, Machine Learning
+Leveraging offline data is a promising way to improve the sample efficiency of online reinforcement learning (RL).
+This paper expands the pool of usable data for offline-to-online RL by leveraging abundant non-curated data that is reward-free, of mixed quality, and collected across multiple embodiments.
+Although learning a world model appears promising for utilizing such data, we find that naive fine-tuning fails to accelerate RL training on many tasks.
+Through careful investigation, we attribute this failure to the distributional shift between offline and online data during fine-tuning.
+To address this issue and effectively use the offline data, we propose two essential techniques: i) experience rehearsal and ii) execution guidance.
+With these modifications, the non-curated offline data substantially improves RL’s sample efficiency.
+Under limited sample budgets, our method achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines across 72 visuomotor tasks spanning 6 embodiments.
+On challenging tasks such as locomotion and robotic manipulation, it outperforms prior methods that utilize offline data by a decent margin.
 1 Introduction
-Figure 1: Aggregate performance on 50 manipulation tasks from Meta-World and 22 locomotion tasks from DMControl with pixel inputs.
-We pre-train one single world model per benchmark on non-curated offline data, which can be fine-tuned to largely improve RL’s sample efficiency.
-With 150k online samples, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 using the same samples and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World).
-Figure 2: Overview of generalist world model pre-training (WPT).
-Facing non-curated offline datasets with reward-free, non-expert, and multi-embodiment data, we train a task and embodiment-agnostic world model.
-The pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
-With the success of scaling laws in vision and language (Brown et al., 2020; He et al., 2022; Kirillov et al., 2023; Touvron et al., 2023), the robotics community has started investigating the use of large-scale offline datasets for robot learning (Brohan et al., 2023b; Walke et al., 2023; Brohan et al., 2023a; O’Neill et al., 2023; Khazatsky et al., 2024; Team et al., 2024).
-Large-scale offline datasets enable the trained agents to learn a wide range of skills as well as help the agent adapt to new tasks with limited samples (Team et al., 2024).
-To leverage offline datasets for decision-making, imitation learning (IL) (Brohan et al., 2023b; Team et al., 2024; Doshi et al., 2024) and offline RL (Levine et al., 2020; Kumar et al., 2020; Fujimoto & Gu, 2021; Kumar et al., 2023) are commonly used.
-However, these methods have limitations in terms of both data and algorithmic aspects.
-IL assumes the availability of expert data from human demonstration or specialist RL agents, which usually involves an expensive data collection procedure.
-Further to this, agents learned by IL are usually sensitive to environmental perturbations such as action delay, lighting conditions, sensor noise, etc (Chae et al., 2022; Zare et al., 2024).
-Offline RL aims to learn agents from offline data but suffers from training instability, especially for pixel-based observations (Kumar et al., 2020; Lu et al., 2023).
-Besides the training stability issue, offline RL requires task-specific datasets, labeled with rewards.
-This poses another challenge for many real-world applications, e.g., robotic manipulation, to utilize offline datasets for new tasks, it is necessary to retrospectively annotate the image-based offline data with rewards, which can be challenging and laborious.
-Furthermore, fine-tuning an offline RL agent, so-called offline-to-online RL (Nair et al., 2020; Lee et al., 2022; Zhao et al., 2022; Yu & Zhang, 2023; Nakamoto et al., 2024), requires additional tricks to stabilize training and prevent performance collapse caused by the distributional shift in the fine-tuning stage.
-In contrast to IL and offline RL, which require curated datasets, many real-world scenarios have data from various sources.
-To this end, we focus on the more realistic setting where the offline dataset includes non-curated mixed-quality data consisting of reward-free and non-expert data, collected by multiple agents with varying embodiments.
-This setting has a minimal requirement for offline data, which significantly expands the pool of usable data.
-The primary research question we aim to answer is:
-What is the best way to leverage non-curated offline data for efficient multi-task & multi-embodiment learning?
-Definition 1.1 (Non-curated data).
-Non-curated data encapsulates all observation-action trajectories, i.e., it can consist of reward-free and non-expert data for any embodiment.
-Existing methods for leveraging offline data (e.g.
-IL and offline RL) typically fail given non-curated data.
-For example, IL would require manually filtering the dataset to select expert trajectories and offline RL would require retrospective reward labeling.
-With such non-curated data, a common strategy is to pre-train a visual encoder (Schwarzer et al., 2021; Nair et al., 2022; Parisi et al., 2022; Xiao et al., 2022; Yang & Nachum, 2021; Shang et al., 2024).
-However, visual pre-training alone fails to fully leverage the rich information that exists in the offline dataset, e.g., dynamics model, informative states, or action prior, etc.
-Facing the non-curated offline dataset, training a world model seems a “natural” choice, as it holds the promise of better generalization and better sample efficiency.
-However, model-based approaches are not commonly used in the offline-to-online setting (Rafailov et al., 2023; Wu et al., 2024).
-This is partially due to the performance margin between model-based and model-free approaches not being significant (Yu et al., 2020; Kidambi et al., 2020) while model-based approaches involve higher training complexity.
-In addition, in the typical offline-to-online RL setting, training an offline agent with model-based approaches can be unstable, especially for pixel inputs (Lu et al., 2023), and strong regularization is required during the online training stage (Rafailov et al., 2023).
-In this paper, we show that with careful design choices, world model pre-training unlocks more useable offline data and better fits the pre-train and fine-tune paradigm used for scaling.
-Our key insights are i) instead of pre-training a policy using model-based offline RL as done in previous work (Yu et al., 2020; Kidambi et al., 2020; Rafailov et al., 2023), learning a world model is stable and scalable, ii) unlike previous methods (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) which learns a world model per task, by simply padding actions to unify the action spaces, a single world model can be pre-trained on the non-curated offline data consisting of data sources from different tasks and different embodiments, iii) via fine-tuning, a pre-trained generalist world model can boost RL’s sample efficiency over a wide range of tasks and embodiments, iv) during fine-tuning, experience rehearsal, and execution guidance can largely improve the performance.
+Leveraging offline data offers a promising way to improve the sample efficiency of Reinforcement Learning (RL).
+Prior work has focused primarily on utilizing curated offline data labeled with rewards [1, 2, 3, 4], which is expensive and laborious to obtain.
+For instance, leveraging offline datasets for new robotic manipulation tasks requires retrospectively annotating image-based data with rewards.
+We instead propose expanding the pool of usable offline data by utilizing abundant non-curated data that is reward-free, of mixed quality, and collected across multiple embodiments.
+This leads to our primary research question:
+How can we effectively leverage non-curated offline data for efficient RL?
+Typical offline-to-online RL [5, 6, 7, 8, 9] methods fail to utilize non-curated offline data due to their assumption of structured data with rewards.
+While pre-training visual encoders [10, 11, 12, 13, 14, 15] is a common approach to utilize non-curated offline datasets, it fails to fully leverage the rich information, such as dynamics models, informative states, and action priors.
+On the other hand, learning world models from offline data appears promising for utilizing the non-curated dataset.
+However, prior work has explored world model training primarily in settings with known rewards [16, 17, 18] or expert demonstrations [19, 20, 21] or focused solely on visual prediction [22, 23].
+Recent approaches [24, 25, 26] have developed novel architectures for world model pre-training using in-the-wild action-free data, but paid limited attention to the fine-tuning process.
+As a result, despite being trained on massive datasets, these methods show only marginal improvements over training-from-scratch baselines.
+Additionally, due to the computational costs of RL experiments, previous work evaluated only on a small set of tasks, leaving the effectiveness of the learned world model unclear on broader tasks.
+In contrast, we extensively evaluate our method on 72 visuomotor control tasks – 12 times more than previous baselines – spanning both locomotion and robotic manipulation, demonstrating consistent improvements over existing approaches.
+Through experiments, we observe that naively fine-tuning a world model fails to improve RL’s sample efficiency on many tasks.
+Through careful investigation, we identify the root cause as a distributional shift between offline and online data during fine-tuning.
+Specifically, when the offline data distribution does not sufficiently cover the data distribution of downstream tasks, the pre-trained world models struggle to benefit policy learning due to this distribution mismatch.
+Especially for tasks requiring hard exploration, even when the pre-trained world model performs well on "promising" state-action distributions, the agent often fails to learn because it cannot effectively reach these promising regions.
+Building on these insights, we propose using offline data in both pre-training and fine-tuning stages, in contrast to previous methods that only consider the offline data for world model pre-training [26, 27, 28].
+By this end, we propose Generalist-to-Specialist Adaptation (GSA), which enables generalist world models pre-trained on non-curated data to master a wide range of downstream tasks.
+GSA has two essential techniques: experience rehearsal and execution guidance.
+Experience rehearsal mitigates distributional shift by retrieving task-relevant trajectories from offline datasets, while execution guidance promotes exploration by steering the agent toward regions where the world model has high confidence.
+Figure 1: Overview of GSA (Generalist-to-Specialist Adaptation).
+Facing non-curated offline data with reward-free, mixed-quality, and multi-embodiment data, we train a task and embodiment-agnostic world model.
+Combined with experience rehearsal and execution guidance, the pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
+Equipped with our proposed techniques, GSA demonstrates strong performance across a diverse set of tasks.
+Specifically, under a limited sample budget (150k samples), GSA achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines (DrQ v2 and Dreamer v3), while matching their performance achieved with larger sample budgets.
+On representative challenging tasks, GSA outperforms baselines that leverage offline data as well as state-of-the-art methods using pre-trained world models by a significant margin.
+Additionally, without any modifications, we show that GSA improves task adaptation, enabling agents to efficiently adapt their skills to new tasks.
+To summarize, our contributions are:
+  • (C1)
+    We propose a more realistic setting for leveraging offline data that consists of reward-free and mixed-quality multi-embodiment data.
+  • (C2)
+    We demonstrate that naive world model fine-tuning fails on many tasks due to distributional shift between pre-training and fine-tuning data.
+  • (C3)
+    We propose two essential techniques, experience rehearsal and execution guidance, to mitigate the distributional gap and encourage exploration.
+  • (C4)
+    We present GSA, which effectively leverages non-curated offline data and significantly outperforms existing approaches across a diverse set of tasks.
+2 Related Work
+In this section, we review methods that leverage offline data, including pre-training in the context of RL and world models. See App. C for more related work and Table 1 for an overview comparison.
+RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve sample efficiency in RL.
+One representative approach is offline RL, which trains agents using offline data without environment interaction.
+These methods typically constrain the distance between learned and behavior policies in different ways [2, 3, 29, 30, 31, 32, 33].
+However, policy performance is highly dependent on dataset quality [34].
+To enable continued improvement, offline-to-online RL methods [5, 6, 7, 9, 17] were developed, which fine-tune policies trained with offline RL by interacting with environments.
+MOTO [17] proposes a model-based offline-to-online RL method with reward-labeled data, and requires model-based value expansion, policy regularization, and controlling epistemic uncertainty, while our method leverages reward-free and multi-embodiment data and requires none of the techniques proposed by MOTO.
+Typical offline-to-online RL face training instability challenges [5, 16].
+To mitigate this issue, RLPD [35] is proposed and demonstrates strong performance by simply concatenating offline and online data, but requires reward-labeled task-specific offline data and does not address multi-embodiment scenarios.
+ExPLORe [36] labels reward-free offline data using approximated upper confidence bounds (UCB) to solve hard exploration tasks, but relies on near-expert data for the target tasks, while we consider a more general setting with non-curated data.
+RL with multi-task offline datasets   Recent work has explored multi-task offline RL [4, 18, 37, 38, 39], but requires known rewards.
+PWM [40] and TDMPC v2 [18] train world models for multi-task RL but are limited to state-based inputs and reward-labeled data.
+To handle unknown rewards, approaches like human labeling [41, 42], inverse RL [43, 44], or generative adversarial imitation learning [45] can be used, though these require human labor or expert demonstrations.
+Yu et al.
+[46] assigns zero rewards to unlabeled data, which introduces additional bias.
+Apart from these, there is a line of work that focuses on representation learning from in-the-wild data [10, 12, 14, 27, 47, 48, 49, 50, 51, 52] but fails to utilize rich information in the dataset, such as dynamics.
+Recent studies [24, 26, 53] explore world model pre-training with action-free data, focusing on world model architecture design to utilize the action-free data.
+However, we demonstrate that naive fine-tuning of pre-trained world models fails on challenging tasks, while our method, incorporating experience rehearsal and execution guidance, significantly improves RL performance across 72 tasks.
 Table 1: Comparison with different policy learning methods that leverage offline data.
-Offline RL Off2On RL RLPD MT IL MT Offline RL WPT (ours)
+Offline RL Off2On RL RLPD MT Offline RL GSA (ours)
 Reward-free offline data
@@
 Training stability
-We extensively evaluate WPT on 72 pixel-based continuous control tasks covering locomotion and manipulation with different action spaces, hard exploration, high dimensions, and complex dynamics.
-Under a limited sample budget (150k samples), WPT outperforms previous SOTA methods by a decent margin.
-Specifically, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 under the same sample budget and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World.).
-For example, WPT enables an agent learning to control an Ant robot to walk forward within 100 trails, while widely used learning-from-scratch baselines need 10-30$\times$ samples.
-We further demonstrate that, without any modifications, WPT also helps task adaptation, where the agent is required to continually adapt its skill to new tasks.
-Our results demonstrate strong performance with a large margin over model-free methods when leveraging non-curated offline data.
-We hope that our results motivate further investigation of model-based approaches for using offline data.
-In summary, our contributions include:
-  • We propose a new and more realistic setting for leveraging offline data where the offline data consists only of reward-free and non-expert multi-embodiment data.
-  • We show that generalist world model pre-training along with a series of careful design choices achieves strong performance in 72 visuomotor tasks spanning 6 embodiments.
-  • To facilitate further research with non-curated offline data, we open-source our datasets, including 60K trajectories and 10M transitions from the DMControl and Meta-World benchmarks.
-2 Related Work
-In this section, we review methods which leverage offline data, including pre-training in the context of RL and the so-called generalist agents.
-See Table 1 for a comparison of what types of data the methods can use.
-RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve the sample efficiency in RL.
-One representative method is Offline RL.
-Offline RL trains agents from offline data without interacting with the environment.
-It usually constrains the distance between the learned policy and behavior policies in different ways (Kumar et al., 2019, 2020; Wu et al., 2019; Kostrikov et al., 2021, 2022; Fujimoto & Gu, 2021; Uchendu et al., 2023).
-However, the performance of the learned policy is highly affected by the quality of offline datasets (Yarats et al., 2022).
-To continue improving the policy, offline-to-online RL (Nair et al., 2020; Zhao et al., 2022; Lee et al., 2022; Yu & Zhang, 2023; Rafailov et al., 2023) was proposed.
-Offline RL and offline-to-online RL commonly suffer from training instability and require additional treatments (Lee et al., 2022; Lu et al., 2023).
-RLPD (Ball et al., 2023) shows that off-policy RL delivers strong performance by directly concatenating the offline data with online data.
-However, RLPD still requires reward-labeled task-specific offline data and hasn’t discussed the multi-embodiment cases.
-Recently, ExPLORe (Li et al., 2023) labels reward-free offline data with approximated upper confidence bound (UCB) of rewards to solve hard exploration tasks.
-However, ExPLORe obtains good performance when the offline data includes near-expert data of the target tasks, while we consider a more general setting with non-curated offline data.
-RL with multi-task offline datasets   Recently, several works have applied offline RL in a multi-task setting (Kalashnikov et al., 2021; Yu et al., 2021; Kumar et al., 2023; Julian et al., 2020; Hansen et al., 2024), but require known rewards.
-Georgiev et al.
-(2024) pre-trains a world model for multi-task RL but is designed for state-based inputs and requires reward-labeled offline data for world model pre-training.
-To lift the assumption of known rewards, human labeling (Cabi et al., 2020; Singh et al., 2019), inverse RL (Ng et al., 2000; Abbeel & Ng, 2004) or generative adversarial imitation learning (Ho & Ermon, 2016) can be applied.
-However, this requires human labor or expert demonstrations.
-Yu et al.
-(2022) directly sets zero rewards for unlabelled data, which introduces additional bias.
-Another stream of work leverages in-the-wild data for RL training.
-Most methods focus on representation learning.
-Stooke et al.
-(2021); Yang & Nachum (2021); Schwarzer et al.
-(2021); Shah & Kumar (2021); Yuan et al.
-(2022); Wang et al.
-(2022); Parisi et al.
-(2022); Sun et al.
-(2023); Ze et al.
-(2023) investigate different representation learning methods and show promising performance boosts during fine-tuning.
-Generalist Agents   RL methods usually perform well on a single task (Vinyals et al., 2019; Andrychowicz et al., 2020), however, this contrasts with humans that can perform multiple tasks well.
-Recent works have proposed generalist agents that master a diverse set of tasks with a single agent (Reed et al., 2022; Brohan et al., 2023b; Team et al., 2024).
-These methods typically resort to scalable models and large datasets and are trained via imitation learning (Brohan et al., 2023b, a; O’Neill et al., 2023; Khazatsky et al., 2024).
-In contrast, we seek to train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
-World models   World models learn to predict future observations or states based on historical information.
-World models have been widely investigated in online model-based RL (Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024).
-Recently, the community has started investigating scaling world models (Ha & Schmidhuber, 2018), for example, Hu et al.
-(2023a); Pearce et al.
-(2024); Wu et al.
-(2025); Agarwal et al.
-(2025) train world model with Diffusion Models or Transformers.
-However, these models are usually trained on demonstration data.
-In contrast, we explore the offline-to-online RL setting – closely fitting the successful pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and non-expert data to increase the amount of available data for pre-training.
 3 Methods
-In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a generalist (i.e.
-multi-task & embodiment) world model, given offline data, which rather importantly, includes reward-free and non-expert data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion.
+In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a multi-task & embodiment world model, given offline data, which rather importantly, includes reward-free, mixed-quality data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion.
 See Fig.
-2 and Alg.
+1 for the overview and Alg.
+1 for the full algorithm.
+3.1 Problem Setup
+In this paper, we assume the agent has access to a non-curated but in-domain offline dataset $\mathcal{D}_{\text{off}}$ with three key characteristics: (i) trajectories lack reward labels $r^{i}_{t}$, (ii) data quality is mixed, and (iii) data comes from multiple embodiments.
+During fine-tuning, the agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=1}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$.
+Our goal is to learn a high-performance policy by leveraging both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ while minimizing the required online interactions $N_{\text{on}}$.
+3.2 Multi-Embodiment World Model Pre-training
+During pre-training, rather than training separate models per task as in previous work [54, 55, 56], we train one world model per benchmark and demonstrate that a single multi-task & embodiment world model can effectively leverage non-curated data.
+Since our primary goal is enabling RL agents to use non-curated offline data rather than proposing a new architecture, we adopt the widely-used recurrent state space model (RSSM) [57] with several modifications: (i) removal of task-related losses, (ii) zero-padding of actions to unify dimensions across embodiments, and (iii) scaling the model to 280M parameters.
+With these changes, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for various tasks.
+Our first stage pre-trains the following components:
+Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{Encoder}:z_{t}\sim q_{\theta}(z_{t}\,|\,h_{t},o_{t})$
+Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}\,|\,h_{t})~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{Decoder}:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}\,|\,h_{t},z_{t}).$
+The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are jointly optimized by minimizing:
+$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})}_{\text{pixel reconstruction loss}}+\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}\,|\,h_{t},o_{t})\,\|\,p_{\theta}(z_{t}\,|\,h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
+The first term minimizes reconstruction error while the second enables latent dynamics learning.
+While there is room to improve world model pre-training through recent self-supervised methods [58] or advanced architectures [59, 60], such improvements are orthogonal to our method and left for future work.
+3.3 RL-based Fine-Tuning with Rehearsal and Guidance
+In our fine-tuning stage, the agent interacts with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$.
+This data is used to learn a reward function via supervised learning while fine-tuning the world model with Eq.
 1.
-3.1 Problem Setup
-In this paper, we assume the agent can access a static offline dataset $\mathcal{D}_{\text{off}}$, which consists of trajectories $\{\tau_{\text{off}}^{i}\}_{i=1}^{N_{\text{off}}}$, where $\tau_{\text{off}}^{i}$ includes observations and actions $\{o^{i}_{t},a^{i}_{t}\}_{t=0}^{T}$ collected by unknown behavior policies.
-This means that (i) rewards $r^{i}_{t}$ are unknown, (ii) $\mathcal{D}_{\text{off}}$ does not necessarily include expert trajectories, and (iii) datasets consist of multi-embodiment data.
-The agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$.
-The goal of this paper is to learn a high-performance policy from both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ whilst minimizing the amount of online interaction ($N_{on}$) by leveraging the unlabeled offline data $\mathcal{D}_{\text{off}}$.
-Removing the requirement of rewards and expert trajectories enables the agent to leverage a large set of diverse datasets.
-3.2 Multi-Embodiment World Model Pre-training
-During pre-training, instead of training one model per task – as done in previous work (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) – we show that a single multi-task & embodiment world model more effectively leverages the offline data.
-Compared to methods (Yang & Nachum, 2021; Schwarzer et al., 2021; Yuan et al., 2022; Parisi et al., 2022; Ze et al., 2023) that only pre-train representations, pre-training a world model creates a general understanding of the environment, which we show can further boost agents’ performance during fine-tuning.
-We adopt a widely used world model in the literature, the recurrent state space model (RSSM) (Hafner et al., 2019) by making several modifications: (i) we remove the task-related losses, (ii) we pad the action with zeros to unify the action dimension of different embodiments, and (iii) we scale the model size to 280M.
-Given these modifications, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for mastering different tasks.
-Our first stage consists of pre-training the following components:
-Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})$
-Encoder $:z_{t}\sim q_{\theta}(z_{t}|h_{t},o_{t})$
-Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}|h_{t})$
-Decoder $:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}|h_{t},z_{t}).$
-The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are optimized jointly by minimizing:
-$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}|z_{t},h_{t})}_{\text{pixel reconstruction loss}}\\ +\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}|h_{t},o_{t})\ ||\ p_{\theta}(z_{t}|h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
-The first term minimizes the reconstruction error while the second term enables the latent dynamics learning.
-Note that we removed the task-related objectives, e.g., reward prediction and continue prediction.
-We note that there is plenty of room to improve the world model’s pre-training, e.g., by leveraging recently developed self-supervised training methods (Eysenbach et al., 2023), or advanced architectures (Vaswani et al., 2017; Gu et al., 2022).
-These improvements are orthogonal to our method and we leave it as future work.
-3.3 RL-based Fine-tuning
-In our fine-tuning stage, we allow the agent to interact with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$.
-The newly collected data is used to learn a reward function via supervised learning whilst also fine-tuning the world model with Eq.
+For simplicity, we represent the concatenation of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$.
+The actor and critic are trained using imagined trajectories $\tilde{\tau}$ generated by rolling out the policy $\pi_{\phi}(a\,|\,s)$ in the world model $p_{\theta}$, starting from initial states $p_{0}(s)$ sampled from the replay buffer.
+The critic $v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})$ learns to approximate the distribution over the $\lambda$-return $V_{t}^{\lambda}$, calculated as:
+$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases}$ (2)
+where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})]$ denotes the critic’s expected value predicted.
+The value function $v_{\phi}$ and actor $\pi_{\phi}$ are updated by maximizing the log likelihood and entropy-regularized $\lambda$-return respectively:
+$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\,|\,s_{t})\right],~{}~{}~{}\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}\,|\,s_{t}]\right)\right].$ (3)
+Figure 2: Visualization of Distribution Mismatch.
+Left: At the early stage of fine-tuning, there is a distribution shift between offline data used for world model pre-training and online data used for RL fine-tuning, which hurts performance.
+Middle: Experience rehearsal mitigates the distributional shift issue.
+Right: Quantitatively, at the early stage of fine-tuning, experience rehearsal reduces the Wasserstein distance between the online data and both the offline and expert data.
+Motivation
+While previous methods typically discard non-curated offline data during fine-tuning [26, 28, 53], we find that relying solely on a pre-trained world model often fails, particularly on hard-exploration tasks.
+To understand why, we analyze the Shelf Place task from Meta-World as an illustrative task by visualizing the distributions of offline data $\mathcal{D}_{\text{off}}$ used for world model pre-training and online data $\mathcal{D}_{\text{on}}$ collected during early RL training in Fig.
+2.
+The t-SNE plot in Fig.
+2 (left) reveals a distribution mismatch between $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$, leading to three key issues: (i) The world model’s accuracy degrades on states visited by the early-stage policy, especially when the offline data distribution is narrow, hurting sample efficiency.
+(ii) For hard exploration tasks, the agent struggles to reach high-reward regions, causing the world model to be fine-tuned on a narrow online data distribution and leading to catastrophic forgetting.
+(iii) The policy update in Eq.
+3 relies on imagined trajectories $\tilde{\tau}=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}\,|\,s_{t})p_{\theta}(s_{t+1}\,|\,s_{t},a_{t})$, where $p_{0}(s)$ is sampled from $\mathcal{D}_{\text{on}}$.
+A narrow $p_{0}(s)$ limits the world model to rollout promising trajectories for policy updates.
+To address these challenges, we introduce two key components: i) experience rehearsal, which mitigates distributional shift by retrieving task-relevant trajectories from non-curated datasets (Fig.
+2 middle, right), and ii) execution guidance, which encourages exploration by steering the agent toward regions where the world model has high confidence.
+Experience Rehearsal
+Prior work like RLPD [35] and ExPLORe [36] has shown that replaying offline data can boost RL training.
+However, these methods use small, well-structured offline datasets.
+In our setting, directly replaying non-curated offline data is infeasible since our datasets are $\sim$100× larger and contain diverse tasks and embodiments.
+We propose retrieving task-relevant trajectories $\mathcal{D}_{\text{retrieved}}=\{\tau^{i}_{\text{retrieved}}\}^{N}_{i=1}$ from the non-curated offline data based on neural feature distance between online samples and offline trajectories.
+This filters out irrelevant trajectories, creating a small dataset of task-relevant data.
+Specifically, we compute:
+$\mathbf{D}=\|\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})\|_{2},$ (4)
+where $\text{e}_{\theta}$ is the encoder learned during world model pre-training, and $o_{\text{on}}$ and $o_{\text{off}}$ are initial observations from trajectories in the online buffer and offline dataset, respectively.
+For efficient search the top-k similar trajectories, we pre-compute key-value pairs mapping trajectory IDs to neural features and use Faiss [61], enabling retrieval in seconds.
+The retrieved data is replayed during fine-tuning, so-called experience rehearsal.
+The retrieved data serves three purposes, as shown in Fig.
 1.
-For simplicity, we represent the concatenate of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$.
-The actor and critic are trained with imagined trajectories $\tilde{\tau}(s,a)$ generated by rolling out the policy $\pi_{\phi}(a|s)$ in the world model $p_{\theta}$ starting from the initial state distribution $p_{0}(s)$, which are samples from replay buffer.
-Following DreamerV3, the critic $v_{\phi}(V_{t}^{\lambda}\mid s_{t})$ learns to approximate the distribution
-over the $\lambda$-return $V_{t}^{\lambda}$ calculated as,
-$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases},$ (2)
-where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\mid s_{t})]$ denotes the expected value of the distribution predicted by the critic. It is trained using maximum likelihood by minimizing
-$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\mid s_{t})\right].$ (3)
-The actor $\pi_{\phi}(a_{t}|s_{t})$ is then updated by maximizing the $\lambda$-return regularized by actor entropy $\mathbf{H}[a_{t}|s_{t}]$:
-$\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}|s_{t}]\right)\right].$ (4)
-In practice, we found that using the pre-trained world model alone fails to work well in many cases, especially for hard-exploration tasks.
-To identify the reasons, we need to rethink the role of each component in policy updates.
-Given the policy update rule in Eq.
-4, the policy update is infected by the distribution of imagined trajectories $\tilde{\tau}(s,a)=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}|s_{t})p_{\theta}(s_{t+1}|s_{t},a_{t})$ and the reward model $r_{\theta}(s_{t},a_{t})$, i.e., the initial state distribution $p_{0}(s)$, the policy $\pi_{\phi}(a|s)$, the dynamics model $p_{\theta}(s_{t+1}|s_{t},a_{t})$, and the reward model $r_{\theta}(s_{t},a_{t})$.
-Intuitively, as shown in Fig.
-3, the policy will learn a good policy if the imagined trajectories $\tilde{\tau}$ contain “promising” data for policy learning and the reward function $r_{\theta}(s_{t},a_{t})$ can correctly label those trajectories.
-The world model pre-training gives a better initialization of $p_{\theta}(s_{t+1}|s_{t},a_{t})$.
-However, when the online data distribution is narrow – as in hard exploration tasks – the initial state distribution $p_{0}(s)$ can also be narrow, which prevents the actor from reaching high-reward states during imagination.
-Also, the reward function $r_{\theta}(s,a)$ trained on the narrow online data distribution will incorrectly assign rewards for transitions generated by the world model pre-trained on a broader distribution.
-That is, the agent fails to improve the policy even if “promising” trajectories exist in $\tilde{\tau}$, since incorrect rewards are assigned.
-Furthermore, fine-tuning the world model $p_{\theta}(s_{t+1}|s_{t},a_{t})$ on the narrow distribution causes catastrophic forgetting.
-In the following, we improve the agents’ performance by reusing non-curated offline data for i) augmenting the initial state distribution $p_{0}(s)$, ii) collecting diverse online trajectories, and iii) improving the reward predictions.
-See Fig.
-3 for an illustrative motivation.
-Retrieval-based Experience Rehearsal
-Figure 3: Motivation of using experience rehearsal and execution guidance.
-During fine-tuning, experience rehearsal augments the initial state $p_{0}(s)$ used for model rollout, which enables the model rollout to start from informative states.
-Execution guidance helps the agent to collect data that is close to the offline data distribution, which enriches the diversity of the collected data and improves the reward model’s accuracy on a broader state distribution.
-Experience replay is a simple yet effective approach to prevent catastrophic forgetting (Khetarpal et al., 2022) as well as improving policy learning (Ball et al., 2023; Li et al., 2023).
-In the RL setting, Ball et al.
-(2023) demonstrates promising results by replaying offline data to improve sample efficiency but requires the offline data to be reward-labeled.
-ExPLORe (Li et al., 2023) reuses single-embodiment but reward-free offline data to encourage exploration by labeling offline data with UCB rewards.
-Unlike the replay scheme used in ExPLORe, where the offline datasets are relatively small and well-structured, directly replaying the non-curated offline data can be challenging.
-Our dataset can be 100$\sim$ 100$\times$ 100 × larger and consist of different tasks and embodiments.
-As such, directly labeling the offline dataset with UCB reward or even replaying the dataset can be infeasible.
-To enable the usage of the non-curated offline data, we propose a simple retrieval-based approach to filter the trajectories that are close to the downstream tasks.
-Specifically, we retrieve a subset of trajectories from the non-curated offline data based on the neural feature distance between the online samples and the trajectories in the offline dataset.
-We calculate the distance as :
-$\mathbf{D}=||\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})||_{2},$ (5)
-where $\text{e}_{\theta}$ is a feature extractor and $o_{\text{on}}$ and $o_{\text{off}}$ are images from the initial observation of trajectories from the online buffer and the offline dataset, respectively.
-In practice, we found that neural representation matters for retrieval and the encoder learned via world model pre-training works better than the general visual model R3M (Nair et al., 2022).
-To efficiently search over the whole dataset, we pre-build the key-value pairs between the trajectories ID and its neural features and use Faiss (Douze et al., 2024) for similarity search.
-Once the key-value database is built, the buffer retrieval procedure only takes seconds.
-The retrieved offline buffer, although reward-free, contains embodiment-specific data and is feasible to use.
-The retrieved offline data can be used to prevent catastrophic forgetting in the world model and to augment the initial state distribution $p_{0}(s)$ used for trajectory imagination.
-Figure 4: WPT enables sample-efficient RL training.
-We include 6 tasks from DMControl and Meta-World with pixel inputs.
-The mean and the corresponding 95% confidence interval across 3 seeds are plotted.
-Full results on 72 tasks are included in App.
-D.
+First, it prevents catastrophic forgetting by continuing to train the world model on relevant pre-training data, particularly important for hard exploration tasks with narrow online data distributions.
+Second, it augments the initial state distribution $p_{0}(s)$ during model rollout, enabling the world model to rollout promising trajectories for policy learning.
+Third, as described below, it enables learning a policy prior for execution guidance.
+Unlike RLPD and ExPLORe, we do not use this data to learn a Q-function, eliminating the need for reward labels.
 Execution Guidance via Prior Actors
-The common practice in RL training is to initialize the replay buffer with a random policy and gradually collect new data by interacting with the environment using the training policy.
-When using a pre-trained world model, it is preferred to start the training from a distribution that is close to the offline data distribution.
-This has three benefits: First, offline data usually contains useful information for policy training, such as near-expert trajectories or a broader state-action distribution.
-Second, the distribution shift between the offline and online data may ruin the pre-trained weights, so bringing the initial distribution close to the offline data is a good strategy to prevent this.
-Thirdly, the reward function can get better predictions for states that are close to the offline data distribution.
-This improves policy learning as the initial states used to generate the imagined trajectories $\tilde{\tau}$ are augmented with offline samples.
-To guide the online data collection such that it remains close to the offline data distribution, we simply train a prior actor $\pi_{\text{bc}}$ via behavior cloning on the retrieved buffer.
-During online data collection, we switch between the policy prior $\pi_{\text{bc}}$ and the RL agent $\pi_{\phi}$ according to a pre-defined schedule.
-Specifically, at the beginning of an episode, we decide whether to use $\pi_{\text{bc}}$ according to the scheduled probability.
-If the $\pi_{\text{bc}}$ is selected to use, we then randomly select the starting time step $t_{\text{bc}}$ and duration $H$ of using $\pi_{\text{bc}}$.
-$\pi_{\phi}$ is then used for the rest of the time steps.
-This procedure is similar to JSRL (Uchendu et al., 2023) but with a couple of differences: i) WPT leverages non-curated offline data while JSRL uses task-specific offline data and ii) WPT shows the superiority of a model-based approach for leveraging offline data while JSRL is a model-free approach.
-As a result, we show that WPT obtains better performance than JSRL on tested tasks.
-The full algorithm is presented in App.
-C.
+Standard RL training initializes the replay buffer with random actions and collects new data through environment interaction using the training policy.
+However, offline data often contains valuable information like near-expert trajectories and diverse state-action coverage that should be utilized during fine-tuning.
+Additionally, distribution shift between offline and online data can degrade pre-trained model weights, making it important to guide the online data collection toward the offline distribution at the early training stage.
+To achieve this, we train a prior policy $\pi_{\text{bc}}$ via behavioral cloning on the retrieved offline data $\mathcal{D}_{\text{retrieved}}$.
+During online data collection, we alternate between this prior policy $\pi_{\text{bc}}$ and the RL policy $\pi_{\phi}$ according to a pre-defined schedule.
+Specifically, at the start of each episode, we probabilistically select whether to use $\pi_{\text{bc}}$.
+If $\pi_{\text{bc}}$ is selected, we randomly choose a starting timestep $t_{\text{bc}}$ and duration $H$ during which $\pi_{\text{bc}}$ is active, with $\pi_{\phi}$ used for the rest timesteps.
+While this approach shares similarities with JSRL [33], our method differs in three key aspects: i) we leverage non-curated rather than task-specific offline data, ii) we demonstrate the benefits of a model-based approach over JSRL’s model-free framework, and iii) we randomly switch between policies mid-episode rather than only using $\pi_{\text{bc}}$ at episode start.
+The complete algorithm and theoretical analysis can be found in App.
+H and App.
+B, respectively.
 4 Experiments
-In this section, we investigate our WPT in two different RL settings.
-In the first setting, we focus on how the proposed WPT helps to improve the sample efficiency of RL training.
-In the second setting, we investigate whether WPT enables fast task adaptation, which is particularly important to empower the agent’s life-long learning ability.
-Lastly, we conduct an ablation study to investigate the role of proposed design choices.
-4.1 Dataset
-Our dataset consists of data from two different benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig.
-2 and Fig.
-9 for the full list.
-For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents (Pathak et al., 2017; Rajeswar et al., 2023), which are trained via curiosity without any task-related information.
-For Meta-World, we collect 50k trajectories spinning 50 tasks by executing the pre-trained RL agents from TDMPCv2 (Hansen et al., 2024) with injected Gaussian noise with $\sigma=[0.1,0.3,0.5,1.0,2.0]$ to mimic imperfect demonstrations.
-In total, the offline dataset consists of 60k trajectories with 10M state-action pairs from 6 different embodiments.
-4.2 WPT Enables Sample-efficient Fine-tuning
-Tasks   We evaluate our method on six selected pixel-based continuous control tasks from DMControl and Meta-World.
-The chosen tasks cover different challenges in RL.
-Except for learning from high-dimensional observations, we include a hard exploration version of Cheetah Run by setting zero rewards below a reward threshold and applying an action penalty.
-We include Walker Run and Quadruped Walk for complex dynamics.
-We further include challenging manipulation tasks from MetaWorld.
-Baselines   In our experiments, we consider four different baselines for leveraging reward-free offline data.
-Note that as these baselines are not designed to handle multi-embodiment offline data, we preprocess the offline data to only contain task-related data for the baselines.
-Although this is an unfair comparison for WPT as baselines eliminate the difficulty raised with the large non-curated dataset, we still outperform these baselines by a large margin on tested tasks.
-The compared baselines are:
-  • R3M pre-trains visual representation aiming to boost policy learning.
-We use a ResNet-18 backbone and train additional policy and value heads via RL.
-We include this baseline to compare with a representative approach for visual pre-training.
-  • UDS suggests to label offline data with zero.
-UDS is an offline RL method, which cannot be directly used for comparison.
-We instead set zero reward for offline data and use RLPD (Ball et al., 2023) for policy training.
-  • ExPLORe labels offline data with UCB reward. We follow the implementation from Li et al. (2023) but use ensembles for reward estimation, which has better performance.
-  • JSRL-BC (Uchendu et al., 2023) collects online data with both the training policy and a prior policy trained on offline data via IL.
-We include JSRL-BC to compare with different ways of using offline data.
-Figure 5: WPT enables fast task adaptation.
-We train an RL agent to control an Ant robot from DMControl to finish a series of tasks incrementally.
-WPT outperforms the widely used baseline by a decent margin by properly leveraging non-curated offline data.
+Figure 3: Left: Quantitative comparison across 72 diverse tasks from Meta-World and DMControl.
+GSA achieves a 102.8% relative improvement in aggregate score over learning-from-scratch baselines when using the same sample budget (150k).
+It also matches the performance of baselines even when they are trained with substantially more samples (see App.
+I for full results).
+Right: Learning curves on representative challenging locomotion and robotic manipulation tasks.
+GSA consistently outperforms state-of-the-art methods that leverage offline data by a decent margin.
+We plot the mean and corresponding 95% confidence interval.
+In the experiments, we aim to answer the following questions: (i) How does GSA compare to state-of-the-art methods that leverage offline data and train-from-scratch baselines in terms of sample efficiency and final performance?
+(ii) How does GSA compare to other leading model-based approaches that utilize offline data?
+(iii) How effectively does GSA adapt to new tasks in a continual learning setting?
+We further conduct detailed ablation studies to thoroughly evaluate our method.
+Tasks
+We evaluate our method on pixel-based continuous control tasks from DMControl and Meta-World.
+The chosen tasks include both locomotion and manipulation tasks covering different challenges in RL, including high-dimensional observations, hard exploration, and complex dynamics.
+We use three random seeds for each task.
+Dataset
+Our dataset consists of data from two benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig.
+12.
+For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents [28, 62], trained via curiosity without task-related information.
+For Meta-World, we collect 50k trajectories across 50 tasks using pre-trained RL agents from TDMPC v2 [18].
+To create a diverse dataset with varying quality, we inject Gaussian noise with $\sigma\in\{0.1,0.3,0.5,1.0,2.0\}$ during execution.
+Combined with the DMControl data, our complete offline dataset contains 60k trajectories (10M state-action pairs) spanning 6 embodiments.
+4.1 GSA Improves Sample Efficiency Across Diverse Tasks
+Comparison with Methods that Leverage Offline Data
+We compare GSA against several state-of-the-art methods that leverage reward-free data to improve RL training: (i) R3M [11], a visual representation pre-training approach that serves as our baseline for comparing pre-trained visual features using non-curated offline data.
+(ii) UDS [46], which assigns zero rewards to offline data.
+Since UDS is an offline RL method, we combine it with RLPD [35] for policy training.
+(iii) ExPLORe [36], which labels offline data using UCB rewards.
+We enhance the original implementation with reward ensembles.
+(iv) JSRL-BC [33], which collects online data using a mixture of the training policy and a behavior-cloned prior policy learned from offline data.
+As the compared baselines cannot handle multi-embodiment data like GSA, we preprocess the offline data to only include task-relevant trajectories for them.
+Despite the baselines having access to better-structured data, GSA still significantly outperforms all baselines across the tested tasks.
+See App.
+G for the details of baselines.
+Results   Fig.
+3 (right) shows comparison results with baselines.
+Our method outperforms all compared baselines by a large margin.
+Compared to R3M, GSA shows the importance of world model pre-training and reusing offline data during fine-tuning, versus representation learning alone.
+R3M fails to improve sample efficiency on most tasks, consistent with findings in Hansen et al.
+[63].
+UDS and ExPLORe reuse offline data by labeling it with zero rewards and UCB rewards, respectively, and concatenating it with online data for off-policy updates.
+UDS shows only slightly better performance on Walker Run compared to R3M and JSRL-BC, demonstrating the ineffectiveness of zero-reward labeling.
+ExPLORe performs better on 2/3 locomotion tasks and shows progress on challenging manipulation tasks, but GSA still significantly outperforms it, demonstrating the superiority of leveraging a generalist world model and properly reusing offline data during fine-tuning.
+GSA also clearly outperforms JSRL-BC.
+JSRL-BS’s performance heavily depends on offline data distribution.
+While JSRL-BC can perform well when a good prior actor can be extracted from offline data, it struggles with non-expert trajectories, showing only marginal improvements over other baselines on Cheetah Run Hard, Assembly, and Shelf Place tasks.
+In contrast, GSA effectively leverages non-expert offline data.
+For example, on Quadruped Walk, GSA benefits from exploratory offline data collected by unsupervised RL, enabling pixel-based control within just 100 trials.
+Comparison with Training-from-Scratch Methods
+We compare GSA with two widely used training-from-scratch baselines: DrQ v2 and Dreamer v3, representing model-free and model-based approaches, respectively.
+Figure 3 (left) and App.
+I show comparison results on 22 locomotion and 50 robotic manipulation tasks with pixel inputs from DMControl and Meta-World benchmarks.
+With 150k online samples, GSA achieves a 102.8% relative improvement in aggregate score compared to DrQ v2 and Dreamer v3, matching their performance obtained with 3.3-6.7$\times$ more samples (500k for DMControl, 1M for Meta-World).
+For example, GSA enables an Ant robot to walk forward within 100 trials, while widely used learning-from-scratch baselines require 10-30$\times$ more samples.
+Furthermore, GSA achieves promising performance on hard exploration tasks where learning-from-scratch baselines fail, such as challenging Meta-World manipulation tasks and hard DMControl tasks.
+Figure 4: Comparison with other world model pre-training methods.
+GSA outperforms state-of-the-art model-based methods without relying on techniques used in iVideoGPT, such as reward shaping and demonstration-based replay buffer initialization.
+Comparison with Other Model-Based Methods
+While most multi-task/multi-embodiment world models focus on visual prediction [22, 23] or imitation learning [19, 20], some works like Seo et al.
+[24]Wu et al.
+[25], and iVideoGPT [26] investigate world model pre-training with in-the-wild data for RL.
+These methods typically focus on designing novel or scalable model architectures to leverage the offline data, but lack mechanisms to better leverage offline data during RL fine-tuning.
+Furthermore, due to the cost of RL training, these methods are usually evaluated on limited task sets, making the effectiveness of the pre-trained world model unclear on diverse tasks.
+Figure 4 compares our method with world model pre-training approaches.
+The baseline results are from the iVideoGPT paper.
+Despite extensive pre-training on diverse manipulation data, iVideoGPT and pre-trained DreamerV3 show only marginal improvements over training-from-scratch baselines.
+In contrast, GSA clearly accelerates RL training by properly leveraging non-curated offline data during both pre-training and fine-tuning.
+Notably, baselines in Fig.
+4 use reward shaping and expert replay buffer pre-filling, while GSA uses none of these tricks yet achieves superior performance.
+This highlights that (i) non-curated offline data contains useful information for RL fine-tuning, and ii) GSA can effectively leverage such data.
+Furthermore, GSA could potentially be combined with iVideoGPT to leverage even more diverse offline data in future work.
+4.2 GSA Enables Fast Task Adaptation
+Figure 5: GSA enables fast task adaptation.
+We train an RL agent to control an Ant robot from DMControl to complete a series of tasks incrementally.
+GSA significantly outperforms the widely used baseline PackNet by properly leveraging non-curated offline data.
+We investigate GSA’s benefits for continual task adaptation, where an agent must incrementally solve a sequence of tasks.
+While similar to continual reinforcement learning (CRL) or life-long RL [64, 65], we use a simplified setting with a limited task set.
+Note that CRL has a broad scope; assumptions and experiment setups vary among methods, making it difficult to set up a fair comparison with other methods.
+Rather than proposing a state-of-the-art CRL method, we aim to demonstrate that GSA offers an effective approach to leverage previous data that also fits the CRL setting.
+Setup & Baselines
+We set our continual multi-task adaptation experiment based on the Ant robot from the DMControl.
+Specifically, the agent sequentially learns stand, walk, run, jump, roll, and roll fast tasks with 300K environment steps per task.
+To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain.
+During training, the agent can access all previous experiences and model weights.
+We compare against a widely used baseline PackNet [66], which iteratively prunes actor parameters while preserving important weights (with larger magnitude) for previous skills.
+For each new task, PackNet fine-tunes the actor model via iterative pruning while randomly reinitializing the critic model since rewards are not shared among tasks.
+Results
+Figure 5 shows GSA significantly outperforms PackNet, enabling adaptation within 100 trials per task.
+With limited samples, PackNet achieves only 20–60% of GSA’s episodic returns.
+We attribute GSA’s superior performance to its ability to leverage the diverse offline data through both world model pre-training and fine-tuning with experience rehearsal, and execution guidance.
+Figure 6: Ablation study on key components.
+“P” represents world model pre-training, “ER” means experience rehearsal, and “G” represents execution guidance.
+The combination of pre-trained generalist world models with retrieval-based experience rehearsal and execution guidance boosts RL performance across diverse tasks.
+4.3 Ablations
+Role of Each Component
+We now analyze each component’s contribution using the same set of tasks from Sec.
+4.1.
+As shown in Fig.
+6, world model pre-training shows promising results when the offline data consists of diverse trajectories, such as data collected by exploratory agents (Walker Run), while it fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks.
+We found that experience rehearsal and execution guidance stabilize training and improve performance on hard exploration tasks like Cheetah Run Hard and challenging manipulation tasks from Meta-World.
+This addresses (i) world model pre-training alone, failing to fully leverage rich state and action information from the non-curated offline data and (ii) distributional shift between offline and online data during fine-tuning hurts the learning.
+The proposed retrieval-based experience rehearsal and execution guidance help utilize offline data and accelerate exploration, which together enable GSA to achieve strong performance on a wide range of tasks.
+Figure 7: Comparison of execution guidance versus uncertainty-based reward labeling.
+GSA demonstrates the effectiveness of using execution guidance over uncertainty-based reward labeling on challenging robotic manipulation tasks.
+Figure 8: Impact of fine-tuning different world model components.
+Comparison with Uncertainty-Aware Reward Function
+To leverage reward-free offline data, ExPLORe [36] proposes to label offline data with uncertainty-based rewards.
+To demonstrate the effectiveness of GSA, we compare it with uncertainty-based rewards.
+Specifically, instead of using execution guidance, we use Optimistic Thompson Sampling (OTS) [67] to label the imagined trajectories via model rollout.
+As shown in Fig.
+7, our method outperforms the variant using OTS on hard exploration tasks, Assembly and Stick Pull, by a large margin, showing the effectiveness of using execution guidance.
+Comparison of Fine-Tuning Different Components
+We now investigate the role of different components in the world model during fine-tuning.
+We use the Quadruped Walk task as a representative task for the investigation.
+As shown in Fig.
+8, encoder and decoder as well as latent dynamics play compatible roles during fine-tuning.
+Fine-tuning the full world model yields the best performance on the tested task.
+The full world model is fine-tuned by default in our experiments.
+5 Conclusion
+We propose GSA, a simple yet efficient approach to leverage ample non-curated offline datasets consisting of reward-free, mixed-quality data collected across multiple embodiments.
+GSA pre-trains a generalist world model on the non-curated data and adapts to downstream tasks via RL.
+We show that naive fine-tuning of world models fails to accelerate RL training due to distributional shift and propose two essential techniques - experience rehearsal and execution guidance - to mitigate this issue.
+Equipped with these techniques, we demonstrate that generalist world models pre-trained on non-curated data are able to boost RL’s sample efficiency across a broader range of locomotion and robotic manipulation tasks.
+We compared GSA against a wide set of baselines, including two widely used training-from-scratch methods, five methods that utilize offline data, and one continual learning method.
+Our GSA consistently delivers strong performance over these baselines.
+Extensive ablation studies reveal the effectiveness of the proposed techniques.
+While promising, GSA can be improved in multiple ways: extending to real-world applications, leveraging in-the-wild offline data, and exploring novel world model architectures.
+Acknowledgments
+We acknowledge CSC – IT Center for Science, Finland, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through CSC.
+We acknowledge the computational resources provided by the Aalto Science-IT project.
+We acknowledge funding from the Research Council of Finland (353138, 362407, 352788, 357301, 339730).
+Aidan Scannell was supported by the Research Council of Finland, Flagship program Finnish Center for Artificial Intelligence (FCAI).
+Appendices
+[…]
+[…]
+Appendix A More Results
+Comparison with Imitation Learning Baseline
+To demonstrate the mix-quality property of the non-curated dataset, we compare GSA with Diffusion Policy, a widely used imitation learning approach by modeling the agent with diffusion models.
+From Fig.
+9, we can see that due to the dataset consisting of non-expert data, the diffusion policy fails to demonstrate satisfactory results, while GSA can effectively utilize the offline data.
+Figure 9: Comparison with Diffusion Policy. GSA can effectively handle non-curated offline data while the imitation learning baseline fails.
+Full Results of Comparison with iVideoGPT
+We compare with other model-based approaches on tasks used in iVideoGPT [26].
+We show that GSA outperforms the baselines without using reward shaping and pre-filling the replay buffer with demonstrations.
+This highlights that although non-curated, the offline data can clearly boost RL training, and GSA can effectively use the information in the data.
+Figure 10: Comparsion with model-based approaches for leveraging offline data.
+Role of Each Component
+We offer inter-quartile mean(IQM) and optimality gap for the ablation study of the role of each proposed component.
+Together with the retrieval-based experience rehearsal and execution guidance, a pre-trained generalist world model boosts RL performance on a wide range of tasks.
 Figure 11: Ablation study on the role of each component.
@@
 Together with the proposed retrieval-based experience rehearsal and execution guidance, world model pre-training boosts RL performance on a wide range of tasks.
-Results   Figure 4 shows comparison results with baselines.
-Our method outperforms or matches all compared baselines by a large margin.
-Compared to R3M, WPT stresses the importance of world model pre-training as well as reusing offline data during fine-tuning compared to representation learning only.
-R3M fails to improve sample efficiency on most of the tested tasks.
-The same conclusion is obtained in Hansen et al.
-(2023).
-UDS and ExPLORe reuse offline data by labeling offline data with zero rewards and UCB rewards respectively.
-The labeled offline data is concatenated with online data for off-policy updates.
-We find UDS only shows slightly better performance on the Walker Run task compared to R3M and JSRL-BC, showing the ineffectiveness of directly labeling offline data with zero rewards.
-ExPLORe shows better performance on 2/3 locomotion tasks compared to other baselines and shows signs of meaningful progress on two challenging manipulation tasks.
-However, WPT still outperforms ExPLORe by a large margin.
-Compared to ExPLORe, WPT demonstrates the superiority of world model pre-training as well as execution guidance.
-Compared to JSRL-BC, our method also demonstrates a clear performance boost.
-JSRL-BC’s performance heavily depends on the distribution of offline datasets.
-It can achieve good performance if a good prior actor can be extracted from the offline data.
-However, in our case, since we aim to lift the assumption of expert trajectories, the pre-trained BC agent usually fails to work well in the target task.
-As a result, JSRL-BC is only slightly better than the rest baselines on the Cheetah Run Hard, Assembly, and Shelf Place tasks while achieving similar results on the rest.
-In contrast, WPT works well on leveraging non-expert offline data.
-For example, the Quadruped Walk shows that WPT can nicely benefit from exploratory offline data collected by unsupervised RL, which enables it to control the Ant robot to walk forward from pixel inputs with only 100 trials.
-In Fig.
-1 and App.
-D, we further compare our method with two widely used learning-from-scratch baselines, Dreamerv3 and DrQv2, on 72 tasks.
-We show that by leveraging the non-curated offline data, WPT can clearly improve the sample efficiency on a wide range of tasks.
-Furthermore, our method achieves promising performance on several hard exploration tasks, while learning-from-scratch baselines fail.
-4.3 WPT Enables Fast Task Adaptation
-In many real-world applications, we want the robot to continuously adapt to new tasks.
-In this section, we investigate whether WPT benefits task adaptation.
-In this setting, the agent is required to incrementally solve a sequence of tasks.
-The setting is very similar to continual reinforcement learning (CRL) or life-long RL (Parisi et al., 2019; Khetarpal et al., 2022) but with a limited set of tasks for simplicity.
-Note that CRL has a broad scope; assumptions and experiment setups vary among methods, which makes it difficult to set up a fair comparison with other methods.
-Instead of proposing a SOTA CRL method, we aim to demonstrate that our method offers a simple yet general recipe to leverage previous data that also fits the CRL setting.
-Setup & baselines   We set our continual multi-task adaption experiment based on the Ant robot from the DeepMind Control Suite.
-Specifically, the agent needs to sequentially solve stand, walk, run, jump, roll, and roll fast tasks.
-We train 300K environment steps for each task.
-To solve one task, the agent accesses all previous experiences as well as the model’s weights trained on previous tasks.
-To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain.
-We compare our method with PackNet (Mallya & Lazebnik, 2018), which is a common baseline for continual learning.
-PackNet iteratively prunes the actor’s parameters by keeping the weights with a larger magnitude while re-initializing the rest.
-By doing so, the actor maintains skills learned previously, which may help the learning of new tasks.
-For each new task, we fine-tune the actor model by iterative pruning while randomly initializing the critic model as rewards are not shared among tasks.
-Results   Figure 5 shows policy performance on incremental task adaption.
-As we can see WPT outperforms baselines by a large margin.
-Compared to PackNet, WPT shows great potential for pre-training of world models with non-curated offline data together with experience rehearsal and actor prior.
-For the tested Ant domain, WPT enables the agent to adapt to new tasks within only 100 trials on all six tasks.
-With the limited sample budget, PackNet only achieves $\sim 20-60\%$ episodic returns of WPT.
-We would argue the diversity of the non-curated offline data contributes to WPT’s supreme performance on task adaptation.
-4.4 Ablations
-In this section, we investigate the role of each component of our method.
-We use the same set of tasks used in Fig.
-4.
-As shown in Fig.
-6, world model pre-training shows promising results when the offline dataset consists of diverse trajectories, such as data collected by exploratory agents, while fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks.
-This is due to i) world model pre-training alone failing to fully leverage the rich action prior in the offline data and ii) the distributional shift between offline and online data hurts the world model fine-tuning.
-The proposed retrieval-based experience rehearsal and execution guidance helps the agent benefit from action prior in the offline data and enriches the online sample diversity, which together enables WPT to achieve strong performance on a wide range of tasks.
-5 Conclusion
-We propose WPT, a simple yet efficient approach to leverage non-curated offline data.
-We show that the generalist world model pre-trained on non-curated data can boost RL training spanning multiple tasks and multiple embodiments.
-Together with retrieval-based experience rehearsal and execution guidance, WPT outperforms baselines on a wide range of tasks, including 22 locomotion tasks and 50 manipulation tasks.
-WPT unlocks ample sources of offline data and our results show that leveraging this non-curated data leads to strong performance.
-Nevertheless, our method can be improved in multiple ways, for example, by extending our methods to real-world applications as well as proposing novel world model architectures.
-Impact Statement
+Appendix B Theoretical Analysis
+In this section, we give a theoretical analysis of the main conclusions in our paper.
+B.1 Proof of the Benefits of Experience Retrieval
+Proposition 1.
+Experience retrieval reduces distribution shift during online fine-tuning, compared to using the full offline dataset directly, in the sense that
+$\mathbb{E}_{s\sim p_{\text{retrieved}},\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right]<\mathbb{E}_{s\sim p_{\text{off}},\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right].$ (5)
+Proof.
+Let $p_{\text{off}}(s)$, $p_{\text{on}}(s)$, and $p_{\text{retrieved}}(s)$ denote the state distributions of the non-curated offline dataset $\mathcal{D}_{\text{off}}$, the online dataset $\mathcal{D}_{\text{on}}$, and the retrieved dataset $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, respectively.
+We simplify the notation as
+$\mathbb{E}_{s\sim p,\ s_{\text{on}}\sim p_{\text{on}}}\left[||s-s_{\text{on}}||_{2}\right]\quad\text{as}\quad\mathbb{E}_{s\sim p}\left[d(s,s_{\text{on}})\right].$
+Since $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, the distribution $p_{\text{off}}(s)$ can be expressed as a mixture distribution:
+$p_{\text{off}}(s)=\alpha\cdot p_{\text{retrieved}}(s)+(1-\alpha)\cdot p_{\text{rest}}(s),$
+where $p_{\text{rest}}(s)$ is the distribution over the remaining offline data, and $\alpha=\frac{|\mathcal{D}_{\text{retrieved}}|}{|\mathcal{D}_{\text{off}}|}$ denotes the fraction of samples in the retrieved dataset.
+The expected total variation for the mixture distribution decomposes as:
+$\mathbb{E}_{s\sim p_{\text{off}}}\left[d(s,s_{\text{on}})\right]=\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right].$ (6)
+Assume that $\mathcal{D}_{\text{retrieved}}$ is constructed by selecting states such that $\|s_{\text{retrieved}}-s_{\text{on}}\|<\epsilon$, for some small $\epsilon>0$.
+Consequently, states in $\mathcal{D}_{\text{rest}}$ satisfy $\|s_{\text{rest}}-s_{\text{on}}\|\geq\epsilon$.
+This construction implies the following bounds:
+$\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]$ $<\epsilon^{\prime},$ (7)
+$\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]$ $\geq\epsilon^{\prime},$ (8)
+for some $\epsilon^{\prime}>0$. Therefore, it follows that
+$\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]>\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right].$
+Substituting into Equation (6) yields:
+$\mathbb{E}_{s\sim p_{\text{off}}}\left[d(s,s_{\text{on}})\right]$ $=\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{rest}}}\left[d(s,s_{\text{on}})\right]$
+$>\alpha\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]+(1-\alpha)\cdot\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right]$
+$=\mathbb{E}_{s\sim p_{\text{retrieved}}}\left[d(s,s_{\text{on}})\right].$
+Thus, the expected total variation between the retrieved data and online data is strictly smaller than that between the full offline data and online data. ∎
+Proposition 2.
+Experience retrieval helps prevent catastrophic forgetting during online fine-tuning.
+Definition 1 (Catastrophic Forgetting due to Data Distribution Shift).
+Catastrophic forgetting occurs when a neural network, after training on a new data distribution, experiences a significant performance drop on previously learned tasks due to the overwriting of representations from earlier distributions, caused by biased parameter updates towards the new distribution.
+Proof.
+Following the previous notations, let $\mathcal{D}_{\text{on}}$ and $\mathcal{D}_{\text{retrieved}}$ denote the online dataset and the retrieved offline dataset, respectively.
+The objective in Eq.
+1 can be written as:
+$\mathcal{L}_{\text{mixed}}(\theta)$ $=\mathcal{L}_{\text{on}}(\theta)+\lambda\cdot\mathcal{L}_{\text{retrieved}}(\theta)$
+$=\mathbb{E}_{p_{\theta},q_{\theta},\ (o,a)\sim\mathcal{D}_{\text{on}}}\left[\sum_{t=1}^{T}-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})+\beta\cdot\mathrm{KL}\left(q_{\theta}(z_{t}\,|\,h_{t},o_{t})\ \|\ p_{\theta}(z_{t}\,|\,h_{t})\right)\right]$
+$+\lambda\cdot\mathbb{E}_{p_{\theta},q_{\theta},\ (o,a)\sim\mathcal{D}_{\text{retrieved}}}\left[\sum_{t=1}^{T}-\ln p_{\theta}(o_{t}\,|\,z_{t},h_{t})+\beta\cdot\mathrm{KL}\left(q_{\theta}(z_{t}\,|\,h_{t},o_{t})\ \|\ p_{\theta}(z_{t}\,|\,h_{t})\right)\right].$
+Assuming the $\lambda$ is a monotonic function of $\alpha=\frac{|\mathcal{D}_{\text{retrieved}}|}{|\mathcal{D}_{\text{off}}|}$ and $\lambda>0$, since $\mathcal{D}_{\text{retrieved}}\subset\mathcal{D}_{\text{off}}$, the term $\mathcal{L}_{\text{retrieved}}(\theta)$ acts as a regularizer during online updates, constraining parameter changes on $\mathcal{D}_{\text{on}}$ in a way that preserves performance on the retrieved offline distribution $p_{\text{retrieved}}$.
+This mitigates the risk of catastrophic forgetting by anchoring the model to previously seen data.
+∎
+B.2 Proof of Improved Performance with Execution Guidance
+Proposition 3 (Performance Improvement via Execution Guidance).
+Let $\pi^{e}$ denote an exploration policy and $\pi^{g}$ a guide policy obtained via imitation learning.
+Assume that $\pi^{g}$ outperforms $\pi^{e}$ at the early stage of training, that is, there exists a constant $\delta>0$ such that for all states $s$ in the support of the state visitation distribution:
+$\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]\geq\delta$ (9)
+Let $\tilde{\pi}$ be an $\alpha$-coupled policy (execution guidance) derived from $\pi^{e}$ and $\pi^{g}$, as defined in Definition 1 of [68], such that:
+$P(a\neq a_{g}|s)\leq\alpha,$ (10)
+where $a\sim\pi^{e}(\cdot|s)$ and $a_{g}\sim\pi^{g}(\cdot|s)$.
+Then, the performance of the $\alpha$-coupled policy $\tilde{\pi}$ exceeds that of the exploration policy $\pi^{e}$ by at least:
+$\eta(\tilde{\pi})-\eta(\pi^{e})\geq\frac{\alpha\cdot\delta}{1-\gamma},$ (11)
+where $\gamma\in[0,1)$ is the discount factor.
+Proof.
+By Lemma 1 in Trust Region Policy Optimization (TRPO) [68], the difference in policy performance can be expressed as:
+$\eta(\tilde{\pi})=\eta(\pi^{e})+\mathbb{E}_{\tau\sim\tilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{\pi^{e}}(s_{t},a_{t})\right]$ (12)
+Define the expected advantage at state $s$ as:
+$\bar{A}_{\pi^{e}}(s)=\mathbb{E}_{a\sim\tilde{\pi}(\cdot|s)}[A_{\pi^{e}}(s,a)]$
+Given the definition of an $\alpha$-coupled policy, we can decompose this expectation:
+$\bar{A}_{\pi^{e}}(s)$ $=(1-\alpha)\cdot\mathbb{E}_{a\sim\pi^{e}(\cdot|s)}[A_{\pi^{e}}(s,a)]+\alpha\cdot\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]$ (13)
+Since $\mathbb{E}_{a\sim\pi^{e}(\cdot|s)}[A_{\pi^{e}}(s,a)]=0$ (by definition of the advantage function), we have:
+$\bar{A}_{\pi^{e}}(s)$ $=\alpha\cdot\mathbb{E}_{a\sim\pi^{g}(\cdot|s)}[A_{\pi^{e}}(s,a)]$
+$\geq\alpha\cdot\delta$ (14)
+The policy performance difference can then be bounded as:
+$\eta(\tilde{\pi})-\eta(\pi^{e})$ $=\mathbb{E}_{\tau\sim\tilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\bar{A}_{\pi^{e}}(s_{t})\right]$
+$=\sum_{s}\rho_{\tilde{\pi}}(s)\bar{A}_{\pi^{e}}(s)$
+$\geq\sum_{s}\rho_{\tilde{\pi}}(s)\cdot\alpha\cdot\delta$
+$=\alpha\cdot\delta\cdot\sum_{s}\rho_{\tilde{\pi}}(s),$ (15)
+where $\rho_{\tilde{\pi}}(s)$ is the discounted state visitation frequency under $\tilde{\pi}$.
+Since $\sum_{s}\rho_{\tilde{\pi}}(s)=\frac{1}{1-\gamma}$ (the sum of discounted state visitation frequencies equals the effective horizon length), we have:
+$\eta(\tilde{\pi})-\eta(\pi^{e})$ $\geq\frac{\alpha\cdot\delta}{1-\gamma}$ (16)
+This establishes that the performance improvement of the $\alpha$-coupled policy $\tilde{\pi}$ over the exploration policy $\pi^{e}$ is at least $\frac{\alpha\cdot\delta}{1-\gamma}$, which is strictly positive when $\alpha>0$, $\delta>0$, and $\gamma<1$.
+∎
+Appendix C More Related Work
+Unsupervised RL   In unsupervised RL, an agent explores the environment based on intrinsic motivations, and the models’ parameters are initialized during this self-motivated exploration stage, aiming for fast downstream task learning [28, 62, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
+Our problem setting differs from unsupervised RL in several ways: i) Unsupervised RL interacts with the environment actively while we leverage static offline datasets, ii) unsupervised RL gives a specific focus on designing different intrinsic rewards, while our setting focuses on improving sample efficiency by leveraging unlabeled datasets.
+Generalist Agents   RL methods usually perform well on a single task [79, 80], however, this contrasts with humans who can perform multiple tasks well.
+Recent works have proposed generalist agents that master a diverse set of tasks with a single agent [81, 82, 83, 84].
+These methods typically resort to scalable models and large datasets and are trained via imitation learning [82, 85, 86, 87].
+In contrast, we train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
+World models   World models learn to predict future observations or states based on historical information.
+World models have been widely investigated in online model-based RL [54, 88, 89, 90, 91].
+Recently, the community has started investigating scaling world models [88], for example, Hu et al.
+[92], Pearce et al.
+[93], Wu et al.
+[26], Agarwal et al.
+[94] train world models with Diffusion Models or Transformers.
+However, these models are usually trained on demonstration data.
+In contrast, we explore the offline-to-online RL setting – closely fitting the pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and multi-embodiment data to increase the amount of available data for pre-training.
+We further identify the distributional shift issue when fine-tuning the pre-trained world model and mitigate the issue by proposing experience rehearsal and execution guidance.
+Appendix D Limitations
+Although demonstrating strong performance on a diverse set of tasks, our method has the following limitations.
+1) The world model architecture used in our paper is the recurrent state space model.
+This model is built upon RNN, which can be limited for scaling.
+This can be mitigated by using a Transformer and a diffusion-based world model.
+However, we note that the main conclusion of this paper should still be valid.
+2) We do not thoroughly discuss the generalization ability of the pre-trained world model.
+With DMControl tasks, our method shows a promising trend in generalizing to unseen tasks.
+However, generalization to new embodiments or novel configurations is still challenging, which requires even diverse training data.
+3) The non-curated offline data used in our paper, although lifting several key assumptions in previous offline-to-online RL, is still in-domain data, i.e., our current method is not able to leverage the vast in-the-wild data.
+A promising direction is to combine in-the-wild data for pre-training as in [26] and the domain-specific “in-house” data (as used in our paper) for post-training.
+4) We only conduct experiments in the simulator.
+Considering the sample efficiency of our proposed method, it could be promising to conduct experiments on real-world applications.
+Appendix E Impact Statement
 This paper contributes to the field of reinforcement learning (RL), with potential applications including robotics and autonomous machines.
@@
 We encourage a broader dialogue on the ethical use of RL technology and its regulation to prevent misuse.
-Appendices
-Appendix A Implementation Details
-A.1 Behavior Cloning
-The Behavior Cloning methods used in both the execution guidance of WPT and JSRL-BC are the same.
-We use a four-layer convolutional neural network (LeCun et al., 1995) with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm (Ba et al., 2016) after all linear layers.
-We list the used encoder and actor architectures for reference.
+Appendix F Compute Resources
+We conduct all experiments on clusters equipped with AMD MI250X GPUs, 64-core AMD EPYC “Trento" CPUs, and 64 GBs DDR4 memory.
+For pre-training, it takes $\sim 48$ GPU hours for 150k steps.
+For fine-tuning, it tasks $\sim 8$ GPU hours per run for 150K environment steps.
+Note that due to AMD GPUs not supporting hardware rendering, the training time should be longer than using Nvidia GPUs.
+To reproduce the GSA’s results in Fig.
+3, it roughly takes 8 h * 72 tasks * 3 seeds = 1728 GPU hours.
+Appendix G Implementation Details
+G.1 Behavior Cloning
+The Behavior Cloning methods used in both the execution guidance of GSA and JSRL-BC are the same.
+We use a four-layer convolutional neural network [95] with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm [96] after all linear layers.
+We list the adopted encoder and actor architectures for reference.
 1class Encoder(nn.Module):
@@
 17 reutrn self.policy(h)
-A.2 JSRL+BC
-Jump-start RL (Uchendu et al., 2023) is proposed as an offline-to-online RL method.
+G.2 JSRL+BC
+Jump-start RL [33] is proposed as an offline-to-online RL method.
 It includes two policies, a prior policy $\pi_{\theta_{1}}(a|s)$ and a behavior policy $\pi_{\theta_{2}}(a|s)$, where the prior policy is trained via offline RL methods and the behavior policy is updated during the online learning stage.
@@
 We then execute the prior policy for $h$ steps and switch to the behavior policy until the end of an episode.
-A.3 ExPLORe
+G.3 ExPLORe
 For the ExPLORe baseline, we follow the original training code   [^1] Source code of ExPLORe https://github.com/facebookresearch/ExPLORe.
@@
 To achieve reasonable performance and eliminate the performance gap caused by implementation-level details, we make the following modifications: i) we replace the RND module with ensembles to calculate uncertainty; ii) the reward function share the latent space with actor and critic.
-Appendix B More Ablation Study
-We include per-task results used in Sec. 4.4 in Fig. 7.
-Figure 7: Ablation study of the role of each component.
-We further compare results using uncertainty-based rewards as suggested in (Li et al., 2023) compared with our expectation guidance.
-Instead of using UCB for reward labels, we use Optimistic Thompson Sampling (OTS) (Hu et al., 2023b), which is less sensitive w.r.t the selection of hyperparameters.
-As shown in Fig.
-8, our method outperforms the variant using OTS on Assembly and Stick Pull by a large margin, showing the effectiveness of using execution guidance.
-Figure 8: Ablation study of using uncertainty-based reward labeling.
-Appendix C Algorithm
+Appendix H Algorithm
 The full algorithm is described in Alg. 1.
-Algorithm 1 Generalist World Model Pre-Training for Efficient RL
-0:  Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$              World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$              Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
+Algorithm 1 Efficient RL by Guiding Generalist World Models with Non-Curated Offline Data
+Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$      World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$      Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
 // Task-Agnostic World Model Pre-Training
@@
           // Update policy and value function
-        Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with policy $\pi_{\phi_{\text{RL}}}$.
+         Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with $\pi_{\phi_{\text{RL}}}$.
          Update policy $\pi_{\phi_{\text{RL}}}$ and value function $v_{\phi}$ with Eq. 3.
@@
 end for
-Appendix D Full Results
+Appendix I Full Results
 In Table 2 and Table 3, we list the success rate of 50 Meta-World benchmark tasks with pixel inputs.
 In Table 4, we list the episodic return of DMControl of 22 tasks.
-We compare WPT at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples.
+We compare GSA at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples.
 We mark the best result with a bold front at 150k samples and use underlining to mark the highest score overall.
-D.1 Meta-World Benchmark
+I.1 Meta-World Benchmark
 Table 2: Success rate of Meta-World benchmark with pixel inputs.
-Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
+Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k GSA @ 150k
 Assembly 0.0 0.0 0.0 0.0 0.2
@@
 Lever Pull 0.73 0.83 0.0 0.33 0.8
-More results in the next page, see Table 3
+More results see Table 3
 Table 3: Success rate of Meta-World benchmark with pixel inputs (Cont.).
-Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
+Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k GSA (ours) @ 150k
 Peg Insert Side 1.0 1.0 0.0 0.27 1.0
@@
 Medium 0.870 0.900 0.130 0.330 0.800
-D.2 DMControl Benchmark
+I.2 DMControl Benchmark
 Table 4: Episodic return of DMControl benchmark with pixel inputs.
-Tasks Dreamer v3 @ 500k DrQ v2 @ 500k Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
+Tasks Dreamer v3 @ 500k DrQ v2 @ 500k Dreamer v3 @ 150k DrQ v2 @ 150k GSA(ours) @ 150k
 CartPole Balance 994.3 992.3 955.8 983.3 995.0
@@
 Medium 606.8 318.70 204.35 166.05 755.0
-Appendix E Hyperparameters
-In this section, we list important hyperparameters used in WPT.
-Table 5: Hyperparameters used in WPT.
+Appendix J Hyperparameters
+In this section, we list important hyperparameters used in GSA.
+Table 5: Hyperparameters used in GSA.
 Hyperparameter Value
@@
 Imagine horizon 16
-Appendix F Task Visualization
+Appendix K Task Visualization
 Figure 12: Visualization of tasks from DMControl and Meta-World used in our paper.